### v1

# mOSCAR: A Large-scale Multilingual and Multimodal Document-level Corpus

Abstract
Multimodal Large Language Models (mLLMs) are trained on a large amount of text-image data. While most mLLMs are trained on caption-like data only, Alayrac et al. (2022) showed that additionally training them on interleaved sequences of text and images can lead to the emergence of in-context learning capabilities. However, the dataset they used, M3W, is not public and is only in English. There have been attempts to reproduce their results but the released datasets are English-only. In contrast, current multilingual and multimodal datasets are either composed of caption-like only or medium-scale or fully private data. This limits mLLM research for the 7,000 other languages spoken in the world. We therefore introduce mOSCAR, to the best of our knowledge the first large-scale multilingual and multimodal document corpus crawled from the web. It covers 163 languages, 315M documents, 214B tokens and 1.2B images. We carefully conduct a set of filtering and evaluation steps to make sure mOSCAR is sufficiently safe, diverse and of good quality. We additionally train two types of multilingual model to prove the benefits of mOSCAR: (1) a model trained on a subset of mOSCAR and captioning data and (2) a model train on captioning data only. The model additionally trained on mOSCAR shows a strong boost in few-shot learning performance across various multilingual image-text tasks and benchmarks, confirming previous findings for English-only mLLMs. The dataset can be accessed here. [^1] https://oscar-project.github.io/documentation/versions/mOSCAR/
1 Introduction
Multimodal large language models (mLLMs) are trained on large amounts of text-image data (Radford et al., 2021; Yu et al., 2022; Li et al., 2023; Wang et al., 2023; OpenAI, 2023; Gemini Team et al., 2023; Chameleon Team, 2024). The main paradigm until recently was to train a model from a large collection of web-crawled images and their captions (Li et al., 2021; Wang et al., 2022; Chen et al., 2023b). Models such as Flamingo (Alayrac et al., 2022) challenged this paradigm by being additionally trained on interleaved sequences of text and images from web documents, showing state-of-the-art results on various tasks and in-context learning capabilities that are not present in models trained on caption-like data only. Additionally, McKinzie et al. (2024) recently proved that including interleaved text-image data during training was necessary to get good few-shot learning performance. However, the datasets used to train mLLMs are either private (Alayrac et al., 2022), monolingual or multilingual but only medium-scale (Srinivasan et al., 2021). Some attempts have been made to reproduce these datasets (Zhu et al., 2023; Laurençon et al., 2023) but the resulting datasets are only available in English.
Few image-text datasets are multilingual and most of them are obtained by translating English caption-like datasets, such as multilingual Conceptual Captions (Sharma et al., 2018), into multiple languages using neural machine translation (NMT) systems (Surís et al., 2022; Maaz et al., 2024). This presents some drawbacks such as some languages still being poorly translated by current state-of-the-art NMT models (Liu et al., 2020; Costa-jussà et al., 2022) and some cultural subtleties inherent in each language not being fully conveyed. Some efforts have been conducted to collect large-scale multilingual image captioning datasets, such as LAION-5B (Schuhmann et al., 2022), but they are limited to caption data too, are relatively noisy and more importantly contain a non-negligible share of “not safe for work” (NSFW) content such as pædopornographic images (Schuhmann et al., 2022).
This motivated us to collect and release the first large-scale multilingual and multimodal document dataset derived from Common Crawl. [^2] https://commoncrawl.org/. The Common Crawl Foundation is a non-profit organization that crawls the web on a monthly basis. Our dataset, multimodal OSCAR (mOSCAR), follows the OSCAR initiative (Ortiz Suárez et al., 2019; Abadji et al., 2021, 2022) and covers 315M documents in 163 languages, 214B tokens and 1.2B images. Figure 1 shows an example of a document, more can be found in Appendix A.3. We carry out extensive filtering to increase its safety and quality. To prove mOSCAR’s utility, we train a multilingual OpenFlamingo (Awadalla et al., 2023) from a Gemma-2B language model (Gemma Team et al., 2024) on a subset of mOSCAR and captioning data from LAION-400M (Schuhmann et al., 2021), recaptioned with BLIP (Li et al., 2022), filtered with CLIP (Radford et al., 2021) and translated with NLLB (Costa-jussà et al., 2022). We compare against a similar model trained on captioning data only and show we obtain a strong boost in few-shot learning, confirming previous findings for English (Alayrac et al., 2022; McKinzie et al., 2024; Laurençon et al., 2024). The dataset and models will be made publicly available.
Figure 1: Example of a French document from mOSCAR.
2 Related Work
Large-scale web-based datasets
Numerous datasets have been created by filtering web-crawled data. These include large-scale text-only datasets (Ortiz Suárez et al., 2019; Raffel et al., 2020; Wenzek et al., 2020; Gao et al., 2020; Abadji et al., 2021; Xue et al., 2021; Laurençon et al., 2022; Abadji et al., 2022; Penedo et al., 2023) and multimodal ones (Sharma et al., 2018; Changpinyo et al., 2021; Jia et al., 2021; Schuhmann et al., 2021, 2022; Byeon et al., 2022; Laurençon et al., 2023; Zhu et al., 2023; Gadre et al., 2024). Even if these datasets are not as high quality as smaller and/or hand-crafted ones, they are now the standard to pretrain foundation models, as it has been shown that training bigger models on more data leads to better downstream performances (Brown et al., 2020; Hoffmann et al., 2022; Touvron et al., 2023a, b).
English image-text datasets
The first open-source image-text datasets were manually created, small-scale and English-only (Ordonez et al., 2011; Lin et al., 2014; Plummer et al., 2015; Krishna et al., 2017). Scaling up these datasets was an appealing solution to overcome limitations of previous image-text models; a few works (Sharma et al., 2018; Changpinyo et al., 2021) proposed to collect millions of image-text pairs from the web before filtering them with well-designed steps. Relaxing the filtering steps enabled the collection of more data and led to large-scale datasets to train image-text foundation models (Radford et al., 2021; Li et al., 2021; Schuhmann et al., 2021, 2022; Byeon et al., 2022). However, these datasets generally contain caption-like image-text pairs only, and it is therefore difficult to observe in-context learning abilities similarly to text-only language models trained on raw documents (Raffel et al., 2020). Alayrac et al. (2022) overcome this issue by training their model directly on documents with interleaved image-text data. While their results are promising, their M3W dataset is English-only and private. Recently, open-source efforts (Zhu et al., 2023; Laurençon et al., 2023) have been made to release a similar dataset but they are still monolingual.
Multilingual image-text datasets
Only a few image-text datasets are available in multiple languages. One of the first focused on collecting Google images from short queries based on word frequencies from Wikipedia pages in 98 languages (Hewitt et al., 2018). Later, Srinivasan et al. (2021) proposed the WIT dataset, an image-text dataset composed of Wikipedia pages. Although of high quality, it is only medium-scale even for high-resource languages and there are fewer than 50k unique images for most languages. Another approach lies in bootstrapping multilingual and multimodal data from a model trained with English-only data (Mohammed et al., 2023). While effective for captioning, it is computationally expensive to implement in practice. Other multilingual image-text datasets exist but focus on captions only and are highly domain-specific (Kosar et al., 2022; Leong et al., 2022).
3 Dataset Creation Pipeline
3.1 Data collection
We collect mOSCAR from the Web ARchive Content (WARC) files of three 2023 Common Crawl dumps, processing them using the FastWARC library (Bevendorff et al., 2021). We remove documents smaller than 500 bytes (50% of the documents), as we find they are usually too small to be considered documents and tend to contain noisy text. We then navigate through the entire Document Object Model (DOM) tree with a depth first search algorithm and ChatNoir library (Bevendorff et al., 2018) to extract nodes of interests corresponding to specific HTML tags.
Following previous work, we extract text from the tags that usually contain the main content of web pages (we refer to them as DOM text nodes), i.e. <p>, <h*>, <title>, <description>, <ul>, <ol>, <aside>, <dl>, <dd>, <dt>. Similarly to (Laurençon et al., 2023), we choose to remove <table> content as most often it is irrelevant and difficult to render. We extract all <img> tags (we refer to them as DOM image nodes). We then remove documents with fewer than 3 text nodes (as they do not contain enough text) and more than 30 image nodes (as we found them to be too noisy).
3.2 Language identification
We identify the language of each document using the state-of-the-art open-LID language detector (Burchell et al., 2023), covering 201 languages. We apply open-LID to each DOM text node and keep the three most probable languages with their respective probabilities. The language of the document is then determined by summing over the probabilities of each language detected for each text segment, weighted by the number of characters in the segment [^3] This is to avoid mis-assigning the language due to the presence of many short, non-informative DOM text nodes in the same language (e.g. “Cookies”, “Subscribe”, “Newsletter” etc.) and because language identification is generally less reliable for short segments. and taking the language with the highest score.
3.3 Text-only filtering
We apply a series of filtering steps to the text content of each document independently of the images, with the aim of discarding poor quality documents and cleaning text as best as possible. We first filter at the text-node level and then at the whole document level, before running near-deduplication to keep unique text nodes within a document and unique documents in the dataset.
Text node filtering
We use a set of heuristics (see Appendix A.2) to extract as much human-generated content as possible while discarding noisy text related to ads and website functions (e.g. “Instagram”, “Facebook”). We then keep DOM text nodes with content over 10 bytes. This step, designed to improve the quality of extracted text, removes on average 55% of text nodes.
Document filtering
We mostly filter “not safe for work” (NSFW) content at the document level. We use an English regular expression to detect adult content, similar to the one used by the Université Toulouse 1 Capitole [^4] https://dsi.ut-capitole.fr/blacklists/index_en.php and remove the entire document if there is a match with any of the DOM text nodes’ contents, removing on average 0.5% of documents (mostly English ones). We acknowledge that there is a high probability that this also discards safe content, e.g. we could remove content from certain communities who use some explicit words in a non-sexual way (Sap et al., 2019). However, we explicitly favour recall over precision to minimise the risk of unsafe content. We additionally remove documents containing fewer than five DOM text nodes and fewer than 300 characters after the previous filtering steps, removing 70.6% of documents.
Deduplication
We conduct several types of per-language deduplication at different levels, as this has been shown to improve training efficiency (Abbas et al., 2023). First, we keep unique documents only by removing exact duplicates at the document level. We also remove exact duplicates of text nodes within the same document (4% of text nodes) and near-duplicate text nodes (1% of text nodes) by computing the Levenshtein ratio (Levenshtein, 1966) between all text nodes within the same document and applying a threshold of 0.95. If near-duplicates are found, we keep the first one in the document. Finally, we conduct per language near-deduplication at the document level with MinHashLSH (Broder, 1997; Gionis et al., 1999) following Smith et al. (2022), removing on average 19% of documents: [^5] With some disparity among languages as we found more duplicates for low- than high-resource languages. we turn documents into hashing vectors, compute min hashes from these vectors and perform Locality Sensitive Hashing to remove duplicates [^6] We performed this using the datasketch python library. (see Appendix A.5 for more details).
3.4 Image-only filtering
We downloaded images from the URLs in DOM image nodes using a modified version of the img2dataset toolkit (Beaumont, 2021) that includes an antivirus scan and follows robots.txt instructions to respect the Robots Exclusion Protocol. We then apply a series of filtering steps, first removing images based on heuristics, and then applying multiple NSFW detection models to remove undesirable content. Finally, we conduct a set of deduplication steps.
Rule-based filters
Similarly to previous works (Schuhmann et al., 2021) and to avoid extracting low-resolution images and favicons, we keep images with a minimum height and width of 150 pixels. We restrict the aspect ratio to be between 3 and 1/3 (to remove banners), we remove images if their URLs contain the words “logo”, “banner”, “button”, “widget”, “icon” or “plugin” or if the image name from the URL matches “twitter”, “facebook” or “rss” (to remove logos). This step removes 13.6% of the URLs. At this stage, we downloaded 2.5B images with an average success rate of 55%.
NSFW detection
We use multiple NSFW automatic models to remove as much unsafe content as possible. We first combine two NSFW detectors: nsfw-detector (Laborde, ), a 5-class classifier with a MobileNet (Howard et al., 2017) backbone fine-tuned on 60GB of annotated data and NudeNet, [^7] https://github.com/vladmandic/nudenet an object detector trained to detect different types of nudity in images. We combined the two models as we found the first to be gender-biased while the second gives a large number of false positives for non-human images. Concretely, we consider an image an NSFW candidate if the sum of the probabilities for the classes ‘porn’ and ‘hentai’ is superior to 0.8 using nsfw-detector. We then tag the image as NSFW if one of the sensitive ‘exposed’ classes of NudeNet gets a probability superior to 0.5. We additionally use Safer by Thorn [^8] https://safer.io/, a private pornography detector, and tag the image as NSFW if the probability of the class ‘pornography’ is superior to 0.8. If a document contains an image with an NSFW tag, we remove the entire document from the dataset, which removes 0.5% of images. We manually inspecting 1,000 images of the remaining data and found no NSFW content. We manually inspected 1,000 images of the removed content and found 63.4% of NSFW images.
CSAM content
Child Sexual Abuse Material (CSAM) is widespread on the internet and is therefore likely to be found in such a large-scale dataset crawled from the web. Removing CSAM is challenging as there is no training data nor open-source detection models available as these could be used in a harmful way. We again rely on Safer, a proprietary 3-class classifier trained to detect CSAM and pornography content from images. We tag the image as CSAM if the probability of the class CSAM is superior to 0.4 to favour recall over precision. As mentioned above, if a document contains an image with a CSAM tag, we remove it from the dataset. This step removes 0.07% of the images.
Deduplication
To avoid memorisation issues often seen in models trained on datasets with many duplicated images (Somepalli et al., 2023; Carlini et al., 2023; Webster et al., 2023; Somepalli et al., 2024), we perform deduplication at the image level. We first remove duplicate images within the same document by URL matching (removing 8.7% of URLs). We then compute a perceptual hash (pHash) for each image using the imagehash library [^9] https://github.com/JohannesBuchner/imagehash and remove images with the same pHash within the same document, keeping only the first occurrence. We also limit the number of times an image can appear in the dataset per-language to 10 using both URL matching and perceptual hashing (this removes 2.5% of images). We do this per-language and not across languages as having the same images in documents from different languages could encourage cross-lingual transfer.
3.5 Data decontamination
LLMs and mLLMs are trained on web-crawled data that can contain the benchmarks they are tested on (Dodge et al., 2021). As they are good at memorizing training data (Carlini et al., 2023), this data contamination is problematic. We therefore discard all images with the same perceptual hash as any of the images from the evaluation benchmarks (and their training sets) we use (see Section 5.1). This step removes on average 126,016 images for high-resource languages (up to 300K images for English), 6,862 images for mid-resource languages and 45 images for low-resource languages.
3.6 Text-image joint filtering
Our aim is to obtain truly multimodal documents where all images are related to at least one of the text nodes in some way [^10] We do not limit ourselves to caption-like relation and instead allow all types of text-image relation. and vice versa. We choose to apply joint text-image filtering to discard images and/or text nodes that are irrelevant to the rest of the document (e.g. the case of ads and website functionalities). To do this, we use NLLB-SIGLIP111111siglip-base-patch16-224 as vision encoder and nllb-distilled-600M as text encoder. (Visheratin, 2023), a multilingual version of SIGLIP (Zhai et al., 2023) trained with the encoder of NLLB (Costa-jussà et al., 2022), which covers all mOSCAR languages. [^12] We use the open-clip (Ilharco et al., 2021) model version and the transformers (Wolf et al., 2020) library. We compute cosine similarity scores between all images and all paragraphs [^13] We refer to paragraph as the text content in a DOM text node. within a same document. To remove irrelevant text nodes or images in a document, we mimic a text-image retrieval task, which means we avoid using arbitrary cosine similarity thresholds for each language and can reduce length biases and those in favour of caption-like paragraphs. For each candidate pair we randomly sample 63 negative images and 63 negative similar-length paragraphs from the same language but other documents. We tag the text node (resp. image) as valid if the cosine similarity of the pair is among the top 8 of the text-to-image (resp. image-to-text) similarity scores computed with the candidate text node (resp. image) and all the negative images (resp. text nodes). This means that we tag the text node (resp. image) as valid if it has a significantly higher score than a score computed with a random image (resp. text) for at least one of the images (resp. text node) in the document. We then discard text nodes and images not tagged as valid (on average 35% of the DOM text nodes and 10% of the images within a document). After this filtering step, we apply additional text-only filters to keep documents superior to 100 bytes.
4 Multimodal Open Super-large Crawled Aggregated coRpus (mOSCAR)
mOSCAR is extracted from three Common Crawl dumps from 2023. Due to computational constraints and in order to extract a maximum number of documents for low-resource languages, we extracted all languages from the first dump only. We removed the 6 most high-resource languages from the second dump and only extracted the languages with fewer than 1M documents for the last dump. Table 1 shows a distribution of the total number of languages and their number of documents.
To avoid data poisoning (Carlini et al., 2024), we release a hash (sha512) with each mOSCAR image.
#documents 10M 5M 1M 500K 200K 50K 10K 5K 1K
#languages 10 15 38 49 58 82 129 142 163
Table 1: Number of languages with at least $N$ documents
(a) Number of images
(b) Number of tokens
(c) Number of images and tokens
Figure 2: Distributions of numbers of tokens and images per document
mOSCAR is composed of 315M documents (214B tokens, 1.2B images) from 163 languages. Figure 2 shows the distribution of images and tokens per document and their joint distribution. As shown in Figure 2(a), the mean and median number of images per document is 2 and 3.80.
4.1 Quality vs Diversity
While improving overall data quality, the filtering steps we applied (see Section 3) necessarily have a negative impact on diversity. We therefore study the trade-off between quality and diversity and compare against previously published, well-used datasets.
4.1.1 Text content
Diversity
By contruction, mOSCAR is diverse in terms of number of languages, so we focus on the diversity of mOSCAR’s English documents and compare against mmc4 (Zhu et al., 2023), OBELICS (Laurençon et al., 2023) and the English subset of WIT (Srinivasan et al., 2021). We compute the Vendi score (Friedman and Dieng, 2023) on a set of SimCSE embeddings (Gao et al., 2021) with a RoBERTa encoder (Liu et al., 2019) to evaluate the content diversity. Since embedding-based diversity metrics target content diversity well but are less relevant for lexical diversity (Tevet and Berant, 2021), we measure lexical diversity via the distinct $n$-gram ratio (Li et al., 2016).
Comparison with other datasets
Figure 3: Perplexity of 100K random documents from different datasets.
Vendi score Dist. $n$-gram ratio
mOSCAR 69.05 ($\pm$ 0.14) 0.472 ($\pm$ 0.002)
mmc4 67.93 ($\pm$ 0.12) 0.494 ($\pm$ 0.002)
OBELICS 58.49 ($\pm$ 0.09) 0.488 ($\pm$ 0.001)
WIT 73.30 ($\pm$ 0.09) 0.530 ($\pm$ 0.001)
Table 2: Average text diversity scores ($\pm$ standard error) of text documents.
For content diversity, we randomly sample 30M documents for mOSCAR, mmc4 and OBELICS and 3M documents for WIT and represent the documents by their SimCSE embedding. We compute the Vendi Score with cosine similarity on a randomly sampled subset of 65,536 documents. Table 2 shows that mOSCAR English content is more diverse than mmc4 and OBELICS but less diverse than WIT. For lexical diversity, we randomly sample 3M documents for mOSCAR, mmc4, OBELICS and WIT and compute the distinct $n$-gram ratio on a subset of 8,192 documents for $n$ from 1 to 4. Table 2 shows that mOSCAR is slightly less lexically diverse than OBELICS and mmc4, while WIT is by far the most diverse.
Quality
To evaluate document quality, we focus on English documents and compute their perplexity using Gemma-2B (Gemma Team et al., 2024). Figure 3 shows the kernel density estimation of the distribution of the perplexity of 100K randomly sampled documents from different datasets: mOSCAR is comparable to mmc4 and WIT, while OBELICS appears to be the of the highest quality. mOSCAR is therefore comparable to other interleaved image-text dataset in terms of quality and diversity of its English subset. It is however more diverse than English-only datasets by its multilingual construction and more than 10 times larger than existing multilingual interleaved image-text datasets such as WIT.
4.1.2 Image diversity
Comparison with other datasets
mOSCAR LAION-400M WIT
55.74 ($\pm$ 0.16) 67.59 ($\pm$ 0.16) 36.14 ($\pm$ 0.08)
(a) Comparison of different datasets.
English All
52.36 ($\pm$ 0.18) 54.78 ($\pm$ 2.29)
(b) mOSCAR (English vs. any language).
Table 3: Average Vendi score ($\pm$ standard error) of images.
We compute the Vendi Score on random samples of images for different datasets, comparing the images from English mOSCAR documents with those from Conceptual Captions (Changpinyo et al., 2021), LAION-400M (Schuhmann et al., 2021) and WIT (Srinivasan et al., 2021). We represent each image by its SigLIP [^14] We use siglip-base-patch16-224. (Zhai et al., 2023) embedding and compute the Vendi score on batches of size 65,536 and a total of 1M images for each dataset. In Table 3(a), we notice that the set of images in mOSCAR documents are more diverse than images from WIT documents but less diverse than LAION-400M.
Multilingual diversity
We also compare the diversity of images from English documents and of images sampled from documents of any language (English included). We use multilingual SigLIP (Chen et al., 2023a) trained on WebLI (Chen et al., 2023b) to compute image embeddings used to get the Vendi score. We again use a batch of size 65,536 and a total of 3M images, and we do not sample multiple images from a same document. For the multilingual setting, we randomly sample 50 languages and an equal number of images for each language to build the batch. As we did not do any image deduplication across languages, we could expect to have less diversity in the multilingual setting. However, Table 3(b) shows that the set of images is on average more diverse when sampled from all documents than from English-only documents. This means that the distribution of images is not exactly the same across languages, potentially due to cultural differences.
5 Training a multilingual multimodal language model
We train a multilingual Flamingo-like model on mOSCAR. As adding captioning data to training data has been shown to improve zero-shot performance (McKinzie et al., 2024), we additionally train on LAION-400M, which we re-captioned using BLIP (Li et al., 2022), filtered with CLIP score (Radford et al., 2021) and translated using distilled NLLB-600M (Visheratin, 2023) following the proportion of languages found in mOSCAR. We use Gemma-2B (Gemma Team et al., 2024) as the underlying language model and we train the model on 35M mOSCAR documents and 70M randomly sampled image-text pairs. We also train a model on 300M image-text pairs as a comparison baseline. We additionally compare with OpenFlamingo-3B-MPT (Awadalla et al., 2023) as the translate-test baseline. The full list of languages for training and the implementation details can be found in Appendix A.5.
5.1 Evaluation setup
We evaluate the models using a broad set of image-text multilingual tasks and benchmarks. We use the IGLUE benchmark (Bugliarello et al., 2022) composed of XVNLI, MaRVL (Liu et al., 2021) to test reasoning, xGQA (Pfeiffer et al., 2022) to test visual question answering capabilities and xFlickr&CO (Young et al., 2014; Karpathy and Fei-Fei, 2015; Yoshikawa et al., 2017) for captioning. We also include Crossmodal-3600 (XM3600) (Thapliyal et al., 2022) and MaXM (Changpinyo et al., 2022) as they cover a broader range of languages. To test to what extent models trained on mOSCAR can perform zero-shot multimodal machine translation (MMT), we also test on Multi30K (Elliott et al., 2016, 2017; Barrault et al., 2018) and CoMMuTE (Futeral et al., 2023). For captioning we compute the CideR (Vedantam et al., 2015) score and we tokenize references and model outputs with the Stanford Core NLP tokenizer for English and Stanza (Qi et al., 2020) tokenizers for other languages. To evaluate Multi30k, we compute BLEU (Papineni et al., 2002) score from Sacrebleu (Post, 2018) with 13a tokenization and default parameters. We use accuracy for CoMMuTE. More details can be found in Appendix A.5.3.
5.2 Results
Tables 4 and 5 show the average results across all languages. Full results are available in Appendix A.6. We notice that the multilingual OpenFlamingo trained additionally on mOSCAR gets better results than the model trained on captioning data only while having seen fewer image-text pairs during training. More importantly, when increasing the number of few-shot examples from 0 to 16, it sees gains of on average +8.19 points on VQA benchmarks and +16.07 CideR points on captioning benchmarks. In contrast, the model trained on text-image pairs only sees gains of +2.82 and +9.08 points respectively. In cross-modal machine translation, the model additionally trained on interleaved data is again far better than the one trained on just captioning data, which is not able to translate the Multi30k benchmark at all. [^15] Most of the time, the model is not able to follow the prompt and only outputs the end of sequence token. Moreover, mOSCAR helps the model to learn to zero-shot disambiguate translations as shown by the improved average score on CoMMuTE (63.75) compared to the model trained on captions only (61.36).
Multilingual Open Flamingo trained on mOSCAR & text-image pairs is also better than OpenFlamingo 3B MPT evaluated on translate test benchmarks. However, we obtain the best results (except for MaXM) by evaluating our multilingual Open Flamingo on the translate-test benchmarks since the underlying language model (Gemma-2B) is far better in English than other languages. We also notice that all models struggle with reasoning classification tasks (MaRVL, XVNLI) where they obtain scores close to random guessing.
xFlickR&CO XM3600 xGQA MaXM MaRVL XVNLI Multi30K CoMMuTE
#shots
Multi. OF 0 19.07 8.73 25.08 19.64 49.77 33.01 22.70 63.75
4 34.32 20.59 31.90 23.90 49.67 36.07 22.79 63.65
full 8 36.77 22.15 33.9 24.41 49.72 37.16 23.21 63.00
16 37.63 22.24 35.71 25.38 49.73 35.36 23.48 62.77
Multi. OF 0 9.57 4.21 8.62 4.01 49.88 33.76 00.00 61.36
4 13.20 9.26 13.45 4.15 49.54 32.04 00.00 61.13
cap. only 8 18.00 10.35 12.82 4.88 49.65 33.71 00.01 60.90
16 19.87 12.07 13.37 4.89 49.79 32.70 00.74 60.25
Table 4: Results averaged over all languages.
xGQA MaXM MaRVL XVNLI
#shots
OF-3B MPT 0 18.34 7.68 49.75 32.73
4 22.97 7.82 49.70 35.82
8 28.57 8.32 49.71 31.29
16 31.82 9.04 49.72 33.29
Multi. OF 0 30.16 10.06 49.93 34.66
4 35.55 9.89 48.99 36.10
full 8 36.78 10.12 50.54 39.69
16 37.75 11.49 49.57 37.97
Table 5: Translate-test results averaged over languages.
Figure 4: Score differences averaged over benchmarks and languages between the model trained on mOSCAR + text-image pairs and the model trained only on text-image pairs.
We additionally compare results at different training steps, defined by the number of images seen during training. Figure 4 shows the difference of averaged scores between the model trained on all data and the model trained only on text-images pairs. We notice that the gap first decreases until 20M images seen and keep increasing over training at all training steps after that. Particularly, the gap is wider for few-shot learning.
6 Conclusion, Limitations and Societal Impacts
We introduce mOSCAR, a large-scale multilingual and multimodal dataset covering 163 languages and composed of 315M documents, 214B tokens and 1.2B images. We show that mOSCAR is of good quality, diverse and can be used to train a multilingual and multimodal LLM. We ensure that mOSCAR is as safe as possible by applying a series of filtering steps to remove NSFW content. We however did not conduct any toxicity analysis or evaluate its biases as this is challenging in a multilingual setting. As it is crawled from the internet, it is possible that mOSCAR reflects biases widespread on it. Nevertheless, by its multilingual nature, mOSCAR is a step towards the inclusion of more languages, cultures and people in accessing mLLMs.
Acknowledgements
This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011014232R1, 2023-AD011014232 and 2023-AD011012254 made by GENCI. It was also partly funded by the last three authors’ chairs in the PRAIRIE institute funded by the French national agency ANR as part of the “Investissements d’avenir” programme under the reference ANR-19- P3IA-0001. We deeply thanks the Jean-Zay support team. We also thanks Filip Šedivý for insightful discussions regarding the removal of CSAM, Thorn for having provided access to their CSAM detector, Zeeshan Khan for discussions regarding the training of the models and Victoria Le Fourner for having manually checked subsamples of NSFW images.
Appendix A Appendix
A.1 mOSCAR languages & statistics
Languages Statistics
Lang. name Code Family Script #documents #images #tokens
Acehnese ace_Latn Austronesian Latin 7,803 32,461 2,889,134
Mesopotamian Arabic acm_Arab Afro-Asiatic Arabic 2,274 10,620 1,047,748
Tunisian Arabic aeb_Arab Afro-Asiatic Arabic 7,640 41,570 2,715,187
Afrikaans afr_Latn Indo-European Latin 54,895 247,774 39,956,585
South Levantine Arabic ajp_Arab Afro-Asiatic Arabic 12,098 87,837 5,167,813
Tosk Albanian als_Latn Indo-European Latin 861,678 2,569,164 452,737,251
Amharic amh_Ethi Afro-Asiatic Ge‘ez 39,588 152,646 35,089,019
North Levantine Arabic apc_Arab Afro-Asiatic Arabic 19,904 128,966 9,560,701
Modern Standard Arabic arb_Arab Afro-Asiatic Arabic 3,936,851 15,126,931 3,401,919,964
Najdi Arabic ars_Arab Afro-Asiatic Arabic 60,229 296,741 43,610,873
Moroccan Arabic ary_Arab Afro-Asiatic Arabic 142,386 698,051 204,723,454
Egyptian Arabic arz_Arab Afro-Asiatic Arabic 835,529 4,054,632 653,626,387
Assamese asm_Beng Indo-European Bengali 3,948 9,210 640,390
Asturian ast_Latn Indo-European Latin 165,745 962,723 37,547,944
Awadhi awa_Deva Indo-European Devanagari 29,324 107,483 4,961,635
Central Aymara ayr_Latn Aymaran Latin 27,384 151,889 5,148,970
South Azerbaijani azb_Arab Turkic Arabic 8,274 38,233 5,256,693
North Azerbaijani azj_Latn Turkic Latin 516,021 1,808,060 257,825,849
Bashkir bak_Cyrl Turkic Cyrillic 4,532 17,174 3,038,766
Bambara bam_Latn Manding Latin 7,674 39,190 1,243,332
Balinese ban_Latn Austronesian Latin 1,886 11,266 542,015
Belarusian bel_Cyrl Indo-European Cyrillic 63,309 287,539 72,976,520
Bemba bem_Latn Atlantic–Congo Latin 1,096 7,479 1,340,471
Bengali ben_Beng Indo-European Bengali 270,406 947,035 35,858,814
Bhojpuri bho_Deva Indo-European Devanagari 6,366 28,131 875,463
Banjar bjn_Latn Austronesian Latin 5,427 27,803 1,898,526
Bosnian bos_Latn Indo-European Latin 1,960,599 7,633,049 1,255,000,505
Buginese bug_Latn Austronesian Latin 3,312 18,648 588,678
Bulgarian bul_Cyrl Indo-European Cyrillic 2,591,998 11,670,028 1,760,971,620
Catalan cat_Latn Indo-European Latin 1,153,864 4,736,634 606,447,390
Cebuano ceb_Latn Austronesian Latin 16,990 91,234 10,748,818
Czech ces_Latn Indo-European Latin 3,918,837 13,291,309 2,823,172,996
Central Kurdish ckb_Arab Indo-European Arabic 36,725 136,566 22,322,689
Crimean Tatar crh_Latn Turkic Latin 6,376 24,124 1,742,727
Welsh cym_Latn Indo-European Latin 40,408 165,897 27,748,345
Danish dan_Latn Indo-European Latin 2,076,298 9,559,600 1,238,277,499
German deu_Latn Indo-European Latin 20,662,696 87,976,200 8,544,986,218
Southwestern Dinka dik_Latn Nilo-Saharan Latin 1,712 6,635 1,319,943
Greek ell_Grek Indo-European Greek 4,916,081 15,209,058 2,923,201,041
English eng_Latn Indo-European Latin 52,215,013 207,904,315 33,570,108,782
Esperanto epo_Latn Artificial Latin 25,157 124,996 28,586,195
Estonian est_Latn Uralic Latin 1,040,368 5,217,366 619,215,048
Basque eus_Latn Isolate Latin 849,043 3,445,539 277,145,498
Faroese fao_Latn Indo-European Latin 15,411 60,340 6,691,327
Fijian fij_Latn Austronesian Latin 1,528 8,776 487,388
Finnish fin_Latn Uralic Latin 2,396,033 10,365,333 1,781,044,864
French fra_Latn Indo-European Latin 20,305,739 78,179,601 14,362,579,829
Friulian fur_Latn Indo-European Latin 37,290 256,456 5,949,600
Nigerian Fulfulde fuv_Latn Atlantic-Congo Latin 1,568 7,124 401,852
West Central Oromo gaz_Latn Afro-Asiatic Latin 4,058 11,763 1,786,093
Scottish Gaelic gla_Latn Indo-European Latin 29,710 153,249 14,605,090
Irish gle_Latn Indo-European Latin 68,858 315,132 47,438,400
Galician glg_Latn Indo-European Latin 518,973 2,381,475 217,063,180
Guarani grn_Latn Tupian Latin 490,945 2,416,633 89,921,114
Gujarati guj_Gujr Indo-European Gujarati 23,062 91,320 3,324,866
Haitian Creole hat_Latn Indo-European Latin 257,745 1,570,699 62,847,106
Hausa hau_Latn Afro-Asiatic Latin 25,364 104,934 13,089,932
Hebrew heb_Hebr Afro-Asiatic Hebrew 1,109,591 4,766,483 893,327,320
Hindi hin_Deva Indo-European Devanagari 579,430 1,830,667 122,558,353
Chhattisgarhi hne_Deva Indo-European Devanagari 1,581 7,263 273,174
Croatian hrv_Latn Indo-European Latin 1,719,617 8,425,510 1,010,674,096
Hungarian hun_Latn Uralic Latin 3,534,506 15,390,083 2,831,715,050
Armenian hye_Armn Indo-European Armenian 339,962 1,141,885 205,635,952
Igbo ibo_Latn Atlantic-Congo Latin 11,529 68,049 8,701,070
Ilocano ilo_Latn Austronesian Latin 78,872 523,195 8,116,113
Indonesian ind_Latn Austronesian Latin 7,016,291 17,324,777 3,981,843,468
Icelandic isl_Latn Indo-European Latin 244,676 1,027,465 137,015,973
Italian ita_Latn Indo-European Latin 12,937,153 47,476,971 8,311,790,842
Javanese jav_Latn Austronesian Latin 24,785 135,583 16,908,805
Japanese jpn_Jpan Japonic Kanji 14,415,292 23,893,768 8,923,348,944
Kabyle kab_Latn Afro-Asiatic Latin 18,508 106,730 4,079,553
Kannada kan_Knda Dravidian Kannada 12,978 42,621 1,442,776
Kashmiri kas_Arab Indo-European Arabic 3,109 11,408 5,731,910
Georgian kat_Geor Kartvelian Georgian 354,436 1,304,281 275,223,026
Kazakh kaz_Cyrl Turkic Cyrillic 252,242 732,648 140,049,214
Halh Mongolian khk_Cyrl Mongolic Cyrillic 124,412 508,217 84,535,241
Khmer khm_Khmr Austroasiatic Kher 24,495 122,243 3,043,925
Kinyarwanda kin_Latn Atlantic-Congo Latin 30,401 172,201 12,049,616
Kyrgyz kir_Cyrl Uralic Cyrillic 53,010 199,713 34,404,281
Northern Kurdish kmr_Latn Indo-European Latin 39,262 164,666 23,834,960
Korean kor_Hang Koreanic Hanja 2,614,089 13,563,283 2,006,080,705
Lao lao_Laoo Kra-Dai Lao 50,611 208,768 31,029,380
Ligurian lij_Latn Indo-European Latin 8,751 56,266 2,958,179
Limburgish lim_Latn Indo-European Latin 189,547 1,076,047 42,534,327
Lingala lin_Latn Atlantic-Congo Latin 24,614 152,132 4,053,459
Lithuanian lit_Latn Indo-European Latin 1,688,811 8,869,443 1,161,476,040
Lombard lmo_Latn Indo-European Latin 30,506 151,855 9,058,614
Latgalian ltg_Latn Indo-European Latin 11,948 61,624 4,148,492
Luxembourgish ltz_Latn Indo-European Latin 44,987 246,346 16,676,872
Ganda lug_Latn Afro-Asiatic Latin 1,878 7,215 789,917
Mizo lus_Latn Sino-Tibetan Latin 7,880 26,817 4,978,472
Standard Latvian lvs_Latn Indo-European Latin 896,243 4,141,648 587,653,855
Magahi mag_Deva Indo-European Devanagari 1,097 3,847 205,763
Malayalam mal_Mlym Dravidian Malayalam 14,140 52,679 1,689,010
Marathi mar_Deva Indo-European Devanagari 50,391 163,868 6,689,250
Minangkabau min_Latn Austronesian Latin 9,341 35,309 1,256,931
Macedonian mkd_Cyrl Indo-European Cyrillic 542,250 1,853,070 307,232,151
Maltese mlt_Latn Afro-Asiatic Latin 120,888 709,242 36,097,957
Maori mri_Latn Austronesian Latin 24,322 130,137 24,957,914
Burmese mya_Mymr Sino-Tibetan Mon 8,144 44,188 539,527
Dutch nld_Latn Indo-European Latin 17,096,727 65,606,013 9,670,041,731
Norwegian Nynorsk nno_Latn Indo-European Latin 199,355 1,012,313 67,799,774
Norwegian Bokmål nob_Latn Indo-European Latin 2,229,702 9,698,128 1,294,178,095
Nepali npi_Deva Indo-European Devanagari 31,239 127,193 3,138,539
Nyanja nya_Latn Atlantic-Congo Latin 12,047 67,192 8,596,769
Occitan oci_Latn Indo-European Latin 164,852 671,881 59,309,549
Odia ory_Orya Indo-European Odia 4,319 15,574 378,635
Pangasinan pag_Latn Austronesian Latin 4,214 32,287 546,071
Eastern Panjabi pan_Guru Indo-European Gurmukhi 11,497 46,168 1,887,991
Papiamento pap_Latn Indo-European Latin 55,224 363,015 10,002,655
Southern Pasto pbt_Arab Indo-European Arabic 32,604 110,807 29,170,322
Western Persian pes_Arab Indo-European Arabic 7,048,946 25,200,571 6,210,479,015
Plateau Malgasy plt_Latn Austronesian Latin 32,521 120,673 29,263,848
Polish pol_Latn Indo-European Latin 14,549,605 60,639,244 11,104,144,109
Portuguese por_Latn Indo-European Latin 8,145,664 26,530,423 4,760,063,083
Dari prs_Arab Indo-European Arabic 515,041 2,589,859 517,053,967
Ayacucho Quechua quy_Latn Quechuan Latin 1,578 11,817 362,690
Romanian ron_Latn Indo-European Latin 5,180,171 17,964,048 3,548,291,261
Rundi run_Latn Atlantic-Congo Latin 20,001 67,096 8,686,054
Russian rus_Cyrl Indo-European Cyrillic 15,913,845 69,542,828 18,909,213,208
Sango sag_Latn Atlantic-Congo Latin 2,124 13,556 454,455
Sicilian scn_Latn Indo-European Latin 73,199 424,362 27,110,743
Sinhala sin_Sinh Indo-European Sinhalese 58,767 221,183 14,270,972
Slovak slk_Latn Indo-European Latin 3,008,599 15,067,234 1,963,804,563
Slovenian slv_Latn Indo-European Latin 1,472,025 7,210,285 935,834,754
Samoan smo_Latn Austronesian Latin 12,346 71,359 14,954,824
Shona sna_Latn Atlantic-Congo Latin 12,698 68,782 6,112,600
Sindhi snd_Arab Indo-European Arabic 21,095 74,289 17,647,825
Somali som_Latn Afro-Asiatic Latin 77,343 301,429 34,554,975
Southern Sotho sot_Latn Atlantic-Congo Latin 7,718 43,146 6,156,450
Spanish spa_Latn Indo-European Latin 22,713,366 78,361,087 14,616,773,475
Sardinian srd_Latn Indo-European Latin 675,539 4,059,493 106,159,957
Serbian srp_Cyrl Indo-European Cyrillic 604,557 2,286,171 401,223,741
Sundanese sun_Latn Austronesian Latin 44,310 236,025 13,627,832
Swedish swe_Latn Indo-European Latin 3,302,730 10,860,518 1,779,284,152
Swahili swh_Latn Atlantic-Congo Latin 137,134 593,418 59,454,896
Silesian szl_Latn Indo-European Latin 23,535 132,459 5,996,972
Tamil tam_Taml Dravidian Tamil 36,196 167,669 4,834,946
Tatar tat_Cyrl Turkic Cyrillic 37,188 143,842 22,831,350
Telugu tel_Telu Dravidian Telugu 22,974 81,033 2,273,772
Tajik tgk_Cyrl Turkic Cyrillic 125,236 417,591 90,503,778
Tagalog tgl_Latn Austronesian Latin 151,437 673,814 97,708,639
Thai tha_Thai Kra-Dai Thai 2,983,837 11,621,786 2,839,211,104
Tigrinya tir_Ethi Afro-Asiatic Ge‘ez 2,657 8,707 1,725,422
Tok Pisin tpi_Latn Indo-European Latin 5,063 35,169 460,853
Turkmen tuk_Latn Turkic Latin 13,024 57,354 9,766,999
Turkish tur_Latn Turkic Latin 4,478,700 12,401,091 2,394,669,068
Twi twi_Latn Atlantic-Congo Latin 3,305 13,634 495,220
Uyghur uig_Arab Turkic Arabic 10,713 41,709 6,785,318
Ukrainian ukr_Cyrl Indo-European Cyrillic 2,721,424 10,929,796 1,928,351,595
Urdu urd_Arab Indo-European Arabic 407,098 1,239,125 242,007,283
Northern Uzbek uzn_Latn Turkic Latin 156,632 798,155 89,022,562
Venetian vec_Latn Indo-European Latin 330,611 1,830,777 71,077,531
Vietnamese vie_Latn Viet-Muong Latin 12,621,521 47,411,488 11,616,191,199
Wolof wol_Latn Atlantic-Congo Latin 4,658 20,380 1,596,432
Xhosa xho_Latn Atlantic-Congo Latin 25,950 142,387 15,809,823
Eastern Yiddish ydd_Hebr Indo-European Hebrew 12,486 57,510 17,369,727
Yoruba yor_Latn Atlantic-Congo Latin 56,700 286,933 32,614,558
Yue Chinese yue_Hant Sino-Tibetan Hant 33,671 203,513 24,172,441
Chinese (Simplified) zho_Hans Sino-Tibetan Hanzi 9,861,262 36,152,754 8,078,842,701
Chinese (Traditional) zho_Hant Sino-Tibetan Hant 3,967,966 16,307,258 2,962,854,441
Standard Malay zsm_Latn Austronesian Latin 1,179,744 5,488,632 432,667,199
Zulu zul_Latn Atlantic-Congo Latin 30,717 156,639 11,345,288
Table 6: Languages & Statistics
A.2 Heuristics to increase the quality of documents
We use a set of heuristics to improve the quality of the documents by discarding some text nodes. We first consider text nodes to be written in Latin scripts if more than 50% of the characters are Latin. In detail, we discard the text node if:
  1. 1.
    It is empty.
  2. 2.
    It contains fewer than 5 bytes for Latin scripts and fewer than 15 bytes for non-Latin scripts.
  3. 3.
    More than 30% of the characters are digits.
  4. 4.
    It contains more than one date.
  5. 5.
    It contains the sequence “lorem ipsum”.
  6. 6.
    The ratio of non-alphabetic characters is superior to 0.33.
  7. 7.
    The symbols ‘{’ or ’‘}’ are in the text.
  8. 8.
    The symbols $\geq$’, $\leq$’, ‘>’ or ‘<’ are more than 2 times in the text.
  9. 9.
    “Follow us”, “javascript”, “copyright” or “©” are in the text.
  10. 10.
    The ratio of capitalized letters is superior to 0.2.
  11. 11.
    The text exactly matches with “comment”, “facebook”, “instagram”, “twitter”, “rss”, “newsletter”, “share” or “follow us”.
  12. 12.
    A character is more than 33% of the total number of characters in the string.
We then also apply some filters to clean the text as much as possible:
  1. 1.
    Remove URLs from all documents.
  2. 2.
[…]
Following previous steps, we keep the text node if it is superior to 5 bytes and we keep the final document if it is superior to 100 bytes.
A.3 Examples of documents
Figure 5: Example of a French document.
Figure 6: Example of a Japanese document.
Figure 7: Example of a Russian document.
Figure 8: Example of an Italian document.
Figure 9: Example of a Khmer document.
Figure 10: Example of an Urdu document.
A.4 Text-Image similarity and DOM Tree
As we rely on the DOM Tree to build the documents and the order of appearance of the nodes could differ from HTML rendering, we attempt to assess to what extent it is a relevant way of constructing a multimodal document. To do so, we rely on the results of the text-image joint filtering step where we compute the ranks of relevant text nodes (resp images) for each image. We plot the distribution of the closest most relevant node for each modality in Figures 11(a) and 11(b). We notice that the most relevant node to either a text node or an image is their closest node in the DOM tree. The cumulative distribution function of the distribution of the closest node reaches 25% for nodes positioned between -5 and 5, which confirms the relevance of using the DOM tree to represent a document.
(a) Relative position in the document of relevant text nodes with respect to images.
(b) Relative position in the document of relevant images with respect to text nodes.
Figure 11: Relative positions of most relevant images and text nodes with respect to the other modality.
A.5 Implementation details
A.5.1 Text deduplication parameters
Following previous work, we near-deduplicate documents using MinHashLSH. We first vectorize the documents using HashingVectorizer from scikit-learn with 2,097,152 features computed on 4-grams and 5-grams within word boundaries. We then compute MinHashes from those vectors with 256 permutations and we finally run Locality Sensitive Hashing with a threshold Jaccard Similarity of 0.8 for finding near-duplicates.
A.5.2 Training implementation details
We train multilingual OpenFlamingo on mOSCAR and multilingual text-image pairs. We use a batch of size 64 for mOSCAR and 128 for captioning data, limiting the number of tokens to 256 for mOSCAR and 32 for captioning data. Similarly to Flamingo and OpenFlamingo, text tokens can only attend to the previous image in the sequence. To increase diversity in the training batch, we randomly reject 2/3 of the documents if they contain only one image. We limit the maximum number of images in a sequence to 8. We randomly sample 8 languages per batch and upsample low-resource languages. We train multilingual OpenFlamingo on 43 languages covering all the languages of the benchmarks we evaluate the models on (see Section A.5.3).
We use Gemma-2B as the underlying language model behind multilingual OpenFlamingo and CLIP ViT-L-14 as the image encoder. We add a cross-attention layer after each decoder layer. Following OpenFlamingo, we add the two special tokens <image> and<|endofchunk|>, whose embeddings were trained. Only the Perceiver Resampler, cross-attention layers and these two embeddings were trained; everything else remained frozen. During training, we apply a factor of 0.2 for the captioning data loss function.
We train the model using the Adam optimizer and a maximum learning rate of 1e-4. We use a constant learning rate scheduler with 1875 warm-up steps. We use 4 accumulation gradient steps to have an effective batch of size 256 for mOSCAR and 512 for captioning data. We train the model on 35M documents and 70M image-text pairs on 8 Nvidia A100 for 120h.
A.5.3 Evaluation details
Metric #examples Languages
xFlickr&CO CideR 2,000 Chinese, English, German, Indonesian, Japanese, Russian, Spanish, Turkish
XM3600 CideR 3,600 Arabic, Czech, Danish, German, Greek, English, Spanish, Farsi,
Finnish, French, Hebrew, Hindi, Croatian, Hungarian, Indonesian, Italian,
Japanese, Korean, Dutch, Norwegian, Poland, Portuguese, Romanian,
Russian, Swedish, Telugu, Thai, Turkish, Ukrainian, Vietnamese, Chinese
xGQA Accuracy 9,666 Bengali, German, English, Indonesian, Korean, Portuguese, Russian, Chinese
MaXM Accuracy $\sim$ 170 English, French, Hindi, Hebrew, Romanian, Thai, Chinese
MaRVL Accuracy $\sim$ 1,150 Indonesian, Swahili, Tamil, Turkish, Chinese
XVNLI Accuracy 1,164 English, Arabic, Spanish, French, Russian
Multi30k BLEU 1,000 French, German, Czech
CoMMuTE Accuracy 310 Czech, French, German
Table 7: Overview of the benchmarks used to evaluate our multilingual OpenFlamingo.
We evaluate on a set of eight benchmarks: xFlickr&CO, XM3600, xGQA, MaXM, MaRVL, XVNLI, Multi30k (Test2016 subset) and CoMMuTE; covering 5 different tasks and 43 languages. Details about the languages, the number of examples and the metric used can be found in Table 7. We used the translate-test [^16] Benchmark automatically translated into English. samples provided by the authors of the benchmarks if available. No translate test samples were provided for MaXM, so we translated the test set using the NLLB-600M distilled model. As no training set was available for MaXM, we use the few-shot examples from xGQA. Since we use Stanza tokenizers, we could not evaluate on all languages from XM3600 as 3 of them were not available. Filipino was also not into the list of mOSCAR languages, so we skip this language during evaluation. The CoMMuTE evaluation set involves choosing between two different translations of a same source text (one correct and one incorrect depending on an image provided to disambiguate the text). We use the lowest perplexity between the two translations as the model’s prediction. We also use Multi30k training set as few-shot examples.
Prompting
Following previous works, the zero-shot setting is composed of two few-shot examples without providing the images. The prompts we use for the different tasks are as follows: [^17] We show the prompts we used with one context example.
For captioning tasks, we use the prompt:
<image>Output:[Caption]<|endofchunk|><image>Output:”,
where [Caption] is replaced by the caption.
For visual question answering tasks, we use the prompt:
<image>Question: [Question] Short Answer: [Answer] <|endofchunk|><image>Question: [Question] Short Answer:”,
where [Question] and [Answer] are replaced by the question and the answer respectively.
For multimodal machine translation tasks, we use the prompt:
<image>Sentence:‘[Caption]’. Translation: [Translation] <|endofchunk|><image>Output:”,
where [Caption] is replaced by the sentence to translate and [Translation] is replaced by its translation.
For MaRVL, we use the prompt:
<image> ‘[Statement]’. True of False? [Answer]<|endofchunk|><image> ‘[Statement]’. True of False?”,
where [Statement] is replaced by the statement and [Answer] by the answer. We also concatenate the left and right image into a single image.
For XVNLI, we use the prompt:
<image> ‘[Statement1]’ - ‘[Statement2]’. entailment, neutral or contradiction? Output: [Answer]<|endofchunk|><image> ‘[Statement1]’ - ‘[Statement2]’. entailment, neutral or contradiction? Output:”,
where [Statement1], [Statement2] and [Answer] are replaced by XVNLI test data.
A.6 Detailed results
De En Es Id Ja Ru Tr Zh
#shots
0 28.87 33.42 16.35 33.26 04.33 13.99 03.68 18.65
Multilingual OF 4 53.99 46.53 35.87 48.85 12.44 31.88 11.35 33.64
mOSCAR + caps. 8 56.05 52.99 37.28 52.51 15.13 32.76 11.69 35.78
16 55.44 56.78 37.72 49.06 17.10 35.43 12.72 36.82
0 16.72 24.57 03.80 10.82 02.82 08.20 02.79 06.82
Multilingual OF 4 21.10 31.05 07.52 09.63 03.84 13.21 07.01 12.20
captions only 8 32.56 35.73 13.35 15.85 05.96 18.13 06.97 15.47
16 29.86 40.57 13.75 23.83 06.92 20.40 07.90 15.73
Table 8: Captioning results (CideR scores) on xFlickr&CO. Bold is best result.
Ar Cs Da De El En Es Fa Fi Fr He
#shots
0 06.18 02.02 11.06 09.12 00.90 43.49 18.01 10.41 01.30 20.61 03.21
Multi. OF 4 21.55 05.93 28.21 21.97 03.38 74.71 36.93 26.27 06.48 39.41 10.34
full 8 23.15 06.23 31.33 24.53 03.47 75.29 37.58 28.84 07.65 43.40 10.83
16 23.80 06.86 31.92 24.92 03.59 76.24 38.39 26.78 07.61 43.26 11.78
0 02.24 00.97 06.42 06.46 03.68 10.02 09.32 04.95 01.14 16.15 00.78
Multi. OF 4 05.36 01.36 13.11 11.82 07.78 35.52 19.96 09.62 01.86 22.48 02.29
Caps only 8 06.76 01.40 15.29 14.39 07.21 37.28 21.90 12.19 02.08 23.27 01.71
16 06.25 02.29 17.96 15.11 07.64 48.03 25.39 09.21 02.10 30.16 02.72
Hi Hr Hu Id It Ja Ko Nl No Pl Pt
#shots
0 02.80 01.47 01.85 09.98 11.15 02.07 01.67 18.97 09.63 04.32 15.49
Multi. OF 4 10.62 07.48 05.51 22.63 27.88 16.87 09.24 42.30 22.35 14.20 29.35
full 8 11.12 07.54 05.91 25.39 29.34 19.35 09.99 46.79 23.54 15.43 32.69
16 12.18 07.71 06.03 23.89 29.17 18.84 09.75 46.95 23.79 15.93 31.98
0 02.29 00.97 03.51 02.98 07.96 01.85 01.05 04.88 05.78 00.92 09.79
Multi. OF 4 04.57 01.72 07.57 06.39 16.23 03.47 04.33 11.26 11.99 01.16 15.93
Caps only 8 05.94 02.17 07.83 09.93 15.40 07.93 05.34 11.87 13.79 01.38 17.50
16 06.36 02.42 09.55 11.77 17.43 10.44 06.03 12.98 14.65 01.28 20.32
Ro Ru Sv Te Th Tr Uk Vi Zh
#shots
0 02.19 06.80 11.45 00.87 08.36 03.11 03.04 21.99 07.19
Multi. OF 4 05.63 20.50 25.59 02.29 21.66 10.87 10.43 38.43 19.34
full 8 06.04 23.11 26.44 03.06 25.19 12.95 10.56 39.75 20.18
16 06.43 21.15 26.51 03.42 25.31 13.57 10.45 40.85 20.46
0 02.24 01.93 04.55 00.67 02.34 02.68 00.80 08.55 02.70
Multi. OF 4 05.35 06.29 15.66 00.77 07.21 05.94 01.76 20.69 07.80
Caps only 8 05.18 07.58 14.01 01.00 06.81 08.90 02.73 23.05 08.99
16 05.06 09.06 20.60 01.18 08.35 10.25 03.47 25.16 11.05
Table 9: Captioning results (CideR scores) on XM3600. Bold is best result.
Bn De En Id Ko Pt Ru Zh
#shots
0 20.99 22.57 32.60 22.45 26.44 23.54 25.31 26.77
Multilingual OF 4 25.72 32.71 38.02 30.74 31.49 31.53 31.75 33.26
mOSCAR + caps. 8 26.87 35.60 39.16 34.25 32.25 34.93 33.59 34.55
16 29.11 37.01 40.64 36.37 34.20 37.15 34.87 36.29
0 10.54 06.51 10.43 07.74 07.50 07.79 08.62 09.84
Multilingual OF 4 12.54 11.90 15.78 13.95 13.70 12.01 12.73 15.03
captions only 8 11.62 11.70 17.29 13.86 12.85 11.60 12.65 15.35
16 09.77 11.86 18.37 13.24 12.48 11.25 11.24 14.33
Translate Test
OF-3B MPT 0 18.64 18.67 - 18.36 17.54 19.21 18.88 17.11
4 23.23 23.40 - 22.95 22.46 23.52 22.41 22.85
8 28.22 29.44 - 28.21 27.67 29.58 28.21 28.63
16 31.31 32.58 - 31.82 31.42 32.74 31.62 31.22
0 30.41 32.1 - 29.35 29.99 31.39 29.06 28.81
Multilingual OF 4 34.89 36.32 - 35.50 35.64 36.84 35.05 34.60
mOSCAR + caps. 8 35.95 37.65 - 36.78 37.14 37.81 36.17 35.98
16 36.78 38.78 - 37.52 37.73 38.68 37.91 36.84
Table 10: VQA results on xGQA. Bold is best result.
En Fr Hi He Ro Th Zh
#shots
0 34.24 15.91 18.08 16.43 13.73 25.75 13.36
Multi. OF 4 36.19 21.21 20.77 19.29 19.01 31.72 19.13
mOSCAR + caps 8 36.96 17.80 20.00 20.36 17.96 33.96 23.83
16 36.19 18.18 21.92 20.71 16.90 41.04 22.74
0 09.73 00.38 07.69 01.43 00.00 05.22 03.61
Multi. OF 4 09.34 02.65 05.00 02.50 00.00 05.60 03.97
captions only 8 09.34 01.89 08.08 05.00 01.06 03.36 05.42
16 08.56 01.14 05.00 08.21 00.35 03.36 07.58
Translate test
OF-3B MPT 0 - 12.50 22.31 00.36 10.92 00.00 00.00
4 - 10.98 25.38 00.36 10.21 00.00 00.00
8 - 10.98 27.31 00.36 11.27 00.00 00.00
16 - 13.26 26.54 01.07 13.38 00.00 00.00
0 - 18.18 28.08 00.00 13.73 00.00 00.36
Multi. OF 4 - 15.91 30.38 00.36 12.68 00.00 00.00
mOSCAR + caps 8 - 15.15 30.77 00.00 14.79 00.00 00.00
16 - 15.91 35.77 00.36 16.90 00.00 00.00
Table 11: VQA results on MaXM. Bold is best result.
Id Sw Ta Tr Zh
#shots
Random chance 50.00 50.00 50.00 50.00 50.00
0 50.00 49.46 49.76 49.83 49.80
Multilingual OF 4 49.91 49.55 49.28 49.83 49.80
mOSCAR + caps 8 50.62 48.65 49.68 49.83 49.80
16 50.18 49.01 49.76 49.83 49.90
0 51.33 49.01 49.52 49.83 49.70
Multilingual OF 4 49.73 49.64 49.19 49.41 49.70
captions only 8 49.91 49.10 49.60 49.75 49.90
16 50.09 49.73 49.60 49.75 49.80
Translate test
OF-3B MPT 0 50.00 49.37 49.76 49.83 49.80
4 50.00 49.64 49.52 49.75 49.60
8 49.82 49.46 49.28 50.08 49.90
16 50.00 49.37 49.44 50.00 49.80
0 49.07 49.79 49.52 50.34 49.60
Multilingual OF 4 49.99 49.79 48.23 49.75 49.76
mOSCAR + caps 8 50.00 48.92 50.64 50.42 48.90
16 49.84 50.00 50.24 48.90 49.75
Table 12: Classification results on MaRVL. Bold is best result.
Ar En Es Fr Ru
#shots
Random chance 33.33 33.33 33.33 33.33 33.33
0 32.90 34.02 31.44 33.85 32.82
Multilingual OF. 4 36.94 36.17 34.19 34.54 38.49
mOSCAR + caps. 8 37.80 39.86 34.97 35.74 37.46
16 34.97 38.92 32.30 34.97 35.65
0 35.48 34.02 33.51 34.45 31.36
Multilingual OF. 4 32.04 31.79 32.73 32.22 31.44
captions only 8 34.02 33.76 32.04 35.57 33.16
16 32.04 32.99 33.76 33.17 31.53
Translate test
OF-3B MPT 0 32.65 - 31.01 31.44 35.82
4 36.25 - 35.82 35.57 35.65
8 31.27 - 31.10 31.10 31.70
16 33.68 - 33.25 32.99 33.25
0 34.88 - 34.88 34.54 34.36
Multilingual OF. 4 36.25 - 36.17 35.91 36.08
mOSCAR + caps. 8 39.60 - 39.52 40.29 39.35
16 37.54 - 37.89 37.46 39.00
Table 13: Classification results on XVNLI. Bold is best result.
Cs De Fr
#shots
0 3.22 28.86 36.01
Multi. OF 4 3.16 28.99 36.22
full 8 3.44 28.76 37.41
16 3.73 29.19 37.53
0 0.00 00.00 00.00
Multi. OF 4 0.00 00.00 00.00
caps. only 8 0.00 00.00 00.03
16 0.00 00.40 01.82
[…]
Cs De Fr
#shots
0 59.09 63.67 68.51
Multi. OF 4 58.77 63.67 68.51
full 8 56.82 63.67 68.51
16 57.79 62.67 67.86
0 58.12 61.67 64.29
Multi. OF 4 59.09 61.00 63.31
caps. only 8 59.09 59.34 64.29
16 58.12 58.67 63.96
[…]

### v2

@@
 We therefore introduce mOSCAR, to the best of our knowledge the first large-scale multilingual and multimodal document corpus crawled from the web.
-It covers 163 languages, 315M documents, 214B tokens and 1.2B images.
+It covers 163 languages, 303M documents, 200B tokens and 1.15B images.
 We carefully conduct a set of filtering and evaluation steps to make sure mOSCAR is sufficiently safe, diverse and of good quality.
-We additionally train two types of multilingual model to prove the benefits of mOSCAR: (1) a model trained on a subset of mOSCAR and captioning data and (2) a model train on captioning data only.
+We additionally train two types of multilingual model to prove the benefits of mOSCAR: (1) a model trained on a subset of mOSCAR and captioning data and (2) a model trained on captioning data only.
 The model additionally trained on mOSCAR shows a strong boost in few-shot learning performance across various multilingual image-text tasks and benchmarks, confirming previous findings for English-only mLLMs.
-The dataset can be accessed here.
+The dataset is released under the Creative Commons CC BY 4.0 license and can be accessed here.
 [^1] https://oscar-project.github.io/documentation/versions/mOSCAR/
+mOSCAR: A Large-scale Multilingual
+and Multimodal Document-level Corpus
+Matthieu Futeralthanks: Correspondence to matthieu.futeral@inria.fr   1,2    Armel Zebaze1,4    Pedro Ortiz Suarez5    Julien Abadji1
+Rémi Lacroix3, 6    Cordelia Schmid1,2    Rachel Bawden1    Benoît Sagot1 1Inria     2Département d’informatique de l’ENS, CNRS, PSL Research University 3Institut du développement et des ressources en informatique scientifique, CNRS 4Sorbonne Université, Paris, France     5Common Crawl Foundation     6Université Paris-Saclay
 1 Introduction
@@
 The Common Crawl Foundation is a non-profit organization that crawls the web on a monthly basis.
-Our dataset, multimodal OSCAR (mOSCAR), follows the OSCAR initiative (Ortiz Suárez et al., 2019; Abadji et al., 2021, 2022) and covers 315M documents in 163 languages, 214B tokens and 1.2B images.
+Our dataset, multimodal OSCAR (mOSCAR), follows the OSCAR initiative (Ortiz Suárez et al., 2019; Abadji et al., 2021, 2022) and covers 303M documents in 163 languages, 200B tokens and 1.15B images.
 Figure 1 shows an example of a document, more can be found in Appendix A.3.
@@
 We compare against a similar model trained on captioning data only and show we obtain a strong boost in few-shot learning, confirming previous findings for English (Alayrac et al., 2022; McKinzie et al., 2024; Laurençon et al., 2024).
-The dataset and models will be made publicly available.
+mOSCAR can be accessed here  [^3] https://huggingface.co/datasets/oscar-corpus/mOSCAR.
 Figure 1: Example of a French document from mOSCAR.
@@
 we turn documents into hashing vectors, compute min hashes from these vectors and perform Locality Sensitive Hashing to remove duplicates [^7] We performed this using the datasketch python library.
-(see Appendix A.5 for more details).
+(see Appendix A.6.1 for more details).
+Toxicity filtering
+Toxic content targeting individuals or groups of people is widespread on the internet and can therefore be found in large-scale web-crawled datasets like mOSCAR without appropriate filtering steps.
+To alleviate this issue, we apply the same method used by Costa-jussà et al.
+(2022) and remove documents from mOSCAR based on the presence of a list of “toxic” words for each language [^8] The list of these words for each language can be found here: https://github.com/facebookresearch/flores/tree/main/toxicity.
+As some words in the list can also be used in a non-toxic way based on the context (e.g.: ‘breast’ in English), we tag the document as toxic and remove it from mOSCAR if it contains at least two distinct words in the list.
+This filtering step removes 0.95% of the documents for very high-resource languages (>5M documents), 2.13% for high-resource languages (<5M, >500K), 0.47% for mid-resource languages (<500K, >50K) and 0.64% for low-resource languages (< 50K).
+When manually analysing 1,000 random documents removed by this filtering step in each of the 2 (high-resource) languages we are native speakers of (English and French), we found 568 documents with toxic content.
+Personal Identifiable Information
+Personal Identifiable Information (PII) can be found in large-scale web-crawled datasets, we therefore conducted an additional filtering step to replace all detected PII by place holder strings using regular expressions (Appendix A.6.2 for more details).
+Concretely, we replaced all detected email addresses, phone numbers, credit card numbers, IP addresses and passport numbers.
+Moreover, it was shown that CommonCrawl contains a non neglictible part of API keys in its content  [^9] https://trufflesecurity.com/blog/[…].
+We therefore scanned the dataset with the trufflehog tool  [^10] https://github.com/trufflesecurity/trufflehog to track down residual API keys that could have passed previous filters.
+We found $\sim$200K positive matches and manually check a random sample of 1K positive matches.
+We found only 2 of them to potentially be API keys, other matches are mainly noisy text nodes not related to PII.
+We removed the 200K text nodes from mOSCAR.
 3.4 Image-only filtering
@@
 We then tag the image as NSFW if one of the sensitive ‘exposed’ classes of NudeNet gets a probability superior to 0.5.
-We additionally use Safer by Thorn [^8] https://safer.io/, a private pornography detector, and tag the image as NSFW if the probability of the class ‘pornography’ is superior to 0.8.
 If a document contains an image with an NSFW tag, we remove the entire document from the dataset, which removes 0.5% of images.
@@
 Removing CSAM is challenging as there is no training data nor open-source detection models available as these could be used in a harmful way.
-We again rely on Safer, a proprietary 3-class classifier trained to detect CSAM and pornography content from images.
+We rely on Thorn’s CSAM classifier [^12] https://safer.io/, a proprietary classifier trained to detect CSAM content in images.
 We tag the image as CSAM if the probability of the class CSAM is superior to 0.4 to favour recall over precision.
@@
 We do this per-language and not across languages as having the same images in documents from different languages could encourage cross-lingual transfer.
+Personal Identification Information
+To protect PII in images, we use a lightweight face detector [^14] https://github.com/Linzaer/Ultra-Light-Fast-Generic-Face-Detector-1MB and apply a threshold of 0.99 to detect faces in the images.
+We apply such a high threshold as we found the model to be biased towards detecting faces with high probability in images without any human.
+For each image in mOSCAR, we distribute the bounding boxes of the detected faces so that users can blur them when downloading the images.
+More details are provided in Appendix A.6.2.
 3.5 Data decontamination
@@
 the case of ads and website functionalities).
-To do this, we use NLLB-SIGLIP111111siglip-base-patch16-224 as vision encoder and nllb-distilled-600M as text encoder.
+To do this, we use NLLB-SIGLIP161616siglip-base-patch16-224 as vision encoder and nllb-distilled-600M as text encoder.
 (Visheratin, 2023), a multilingual version of SIGLIP (Zhai et al., 2023) trained with the encoder of NLLB (Costa-jussà et al., 2022), which covers all mOSCAR languages.
@@
 After this filtering step, we apply additional text-only filters to keep documents superior to 100 bytes.
+We also reapply the open-lid language detector (Burchell et al., 2023) as described in Section 3.2 as we found the last filtering step to change the major language of some documents.
 4 Multimodal Open Super-large Crawled Aggregated coRpus (mOSCAR)
@@
 To avoid data poisoning (Carlini et al., 2024), we release a hash (sha512) with each mOSCAR image.
-#documents 10M 5M 1M 500K 200K 50K 10K 5K 1K
-#languages 10 15 38 49 58 82 129 142 163
+mOSCAR is composed of 303M documents (200B tokens, 1.15B images) from 163 languages.
+Figure 2 shows the distribution of images and tokens per document and their joint distribution.
+As shown in Figure 2(a), the mean and median number of images per document is 2 and 3.80.
+#docs. 10M 5M 1M 500K 200K 50K 10K 5K 1K
+#langs. 10 15 36 46 56 75 119 133 154
 Table 1: Number of languages with at least $N$ documents
@@
 (c) Number of images and tokens
-Figure 2: Distributions of numbers of tokens and images per document
-mOSCAR is composed of 315M documents (214B tokens, 1.2B images) from 163 languages.
-Figure 2 shows the distribution of images and tokens per document and their joint distribution.
-As shown in Figure 2(a), the mean and median number of images per document is 2 and 3.80.
+Figure 2: Distributions of numbers of tokens and images per document.
 4.1 Quality vs Diversity
@@
 Since embedding-based diversity metrics target content diversity well but are less relevant for lexical diversity (Tevet and Berant, 2021), we measure lexical diversity via the distinct $n$-gram ratio (Li et al., 2016).
+An analysis of the topics (Grootendorst, 2022) found in multiple languages of mOSCAR where we show diverse topics across languages can be found in Appendix A.2.
 Comparison with other datasets
@@
 55.74 ($\pm$ 0.16) 67.59 ($\pm$ 0.16) 36.14 ($\pm$ 0.08)
-(a) Comparison of different datasets.
-English All
-52.36 ($\pm$ 0.18) 54.78 ($\pm$ 2.29)
-(b) mOSCAR (English vs. any language).
-Table 3: Average Vendi score ($\pm$ standard error) of images.
+Table 3: Average Vendi score ($\pm$ standard error) of images sampled from different datasets.
 We compute the Vendi Score on random samples of images for different datasets, comparing the images from English mOSCAR documents with those from Conceptual Captions (Changpinyo et al., 2021), LAION-400M (Schuhmann et al., 2021) and WIT (Srinivasan et al., 2021).
@@
 (Zhai et al., 2023) embedding and compute the Vendi score on batches of size 65,536 and a total of 1M images for each dataset.
-In Table 3(a), we notice that the set of images in mOSCAR documents are more diverse than images from WIT documents but less diverse than LAION-400M.
+In Table 3, we notice that the set of images in mOSCAR documents are more diverse than images from WIT documents but less diverse than LAION-400M.
+English All
+52.36 ($\pm$ 0.18) 54.78 ($\pm$ 2.29)
+Table 4: Average Vendi score ($\pm$ standard error) of images sampled from mOSCAR (English vs. any language).
 Multilingual diversity
@@
 As we did not do any image deduplication across languages, we could expect to have less diversity in the multilingual setting.
-However, Table 3(b) shows that the set of images is on average more diverse when sampled from all documents than from English-only documents.
+However, Table 4 shows that the set of images is on average more diverse when sampled from all documents than from English-only documents.
 This means that the distribution of images is not exactly the same across languages, potentially due to cultural differences.
+xFlickR&CO XM3600 xGQA MaXM MaRVL XVNLI Multi30K CoMMuTE
+#shots
+Multi. OF 0 16.91 7.45 26.95 22.33 49.56 33.88 22.91 63.34
+4 34.80 22.18 32.33 26.33 49.64 34.07 23.27 63.22
+mOSCAR + cap. 8 36.90 23.48 34.24 27.08 51.48 36.60 23.59 63.54
+16 39.46 23.67 35.23 27.47 49.84 34.85 23.85 62.78
+Multi. OF 0 9.57 4.21 8.62 4.01 49.88 33.76 00.00 61.36
+4 13.20 9.26 13.45 4.15 49.54 32.04 00.00 61.13
+cap. only 8 18.00 10.35 12.82 4.88 49.65 33.71 00.01 60.90
+16 19.87 12.07 13.37 4.89 49.79 32.70 00.74 60.25
+Table 5: Results averaged over all languages.
+Multi.
+OF refers to multilingual Open Flamingo, mOSCAR + cap.
+refers to the model trained on text-image pairs and mOSCAR while cap.
+only refers to the model trained only on text-image pairs.
+Bold is best result.
+xFlickR&CO XM3600 xGQA MaXM MaRVL XVNLI Multi30K CoMMuTE
+#shots
+Multi. OF (35M) 0 19.07 8.73 25.08 19.64 49.77 33.01 22.70 63.75
+4 34.32 20.59 31.90 23.90 49.67 36.07 22.79 63.65
+mOSCAR + cap. 8 36.77 22.15 33.9 24.41 49.72 37.16 23.21 63.00
+16 37.63 22.24 35.71 25.38 49.73 35.36 23.48 62.77
+Multi. OF (35M) 0 9.39 4.67 19.81 14.63 49.71 32.78 26.99 56.75
+4 7.68 2.99 25.68 16.12 49.72 33.51 26.99 53.27
+WIT + cap. 8 8.91 3.63 27.06 16.81 49.74 32.77 26.99 55.33
+16 9.74 4.14 28.14 16.34 49.74 33.63 26.99 54.04
+Table 6: Results averaged over all languages and comparison between a model trained on WIT and a checkpoint of multilingual Open Flamingo trained on 35M mOSCAR documents (full model was trained on 50M mOSCAR documents).
+Both models were trained on 35M documents from their respective training datasets and 70M text-image pairs for fair comparison.
+Multi.
+OF (35M) refers to multilingual Open Flamingo trained on 35M documents.
+Bold is best result.
 5 Training a multilingual multimodal language model
-We train a multilingual Flamingo-like model on mOSCAR.
+We train a multilingual Flamingo-like model on mOSCAR that we call multilingual Open Flamingo.
 As adding captioning data to training data has been shown to improve zero-shot performance (McKinzie et al., 2024), we additionally train on LAION-400M, which we re-captioned using BLIP (Li et al., 2022), filtered with CLIP score (Radford et al., 2021) and translated using distilled NLLB-600M (Visheratin, 2023) following the proportion of languages found in mOSCAR.
-We use Gemma-2B (Gemma Team et al., 2024) as the underlying language model and we train the model on 35M mOSCAR documents and 70M randomly sampled image-text pairs.
-We also train a model on 300M image-text pairs as a comparison baseline.
+We use Gemma-2B (Gemma Team et al., 2024) as the underlying language model and we train the model on 50M mOSCAR documents and 100M randomly sampled image-text pairs.
+We also train a model on 300M image-text pairs, a model trained on 35M WIT (Srinivasan et al., 2021) documents and 70M text-image pairs and a model trained on 50M mOSCAR documents from the English subset and 100M English image-text pairs as comparison baselines.
 We additionally compare with OpenFlamingo-3B-MPT (Awadalla et al., 2023) as the translate-test baseline.
 The full list of languages for training and the implementation details can be found in Appendix A.6.
+#shots xFlickR&CO XM3600 xGQA MaXM XVNLI
+Multilingual OF 0 29.64 42.57 34.24 36.58 34.62
+4 51.47 77.98 37.91 38.13 33.59
+mOSCAR + cap. 8 56.75 77.64 39.44 38.52 38.75
+16 59.89 78.18 40.09 35.80 36.60
+English OF 0 32.70 43.75 34.71 36.19 35.82
+4 51.39 75.33 37.48 37.96 34.88
+English mOSCAR + English cap. 8 51.44 77.73 39.64 38.35 36.86
+16 59.24 78.38 40.36 37.35 37.11
+Table 7: Results on the English subsets of the test sets and comparison between multilingual Open Flamingo and an Open Flamingo trained on the English subset of mOSCAR and English text-image pairs (English OF).
+Both models were trained on 50M documents from their respective training datasets and 100M text-image pairs for fair comparison.
+Bold is best result.
 5.1 Evaluation setup
@@
 More details can be found in Appendix A.6.4.
-5.2 Results
-Tables 4 and 5 show the average results across all languages.
-Full results are available in Appendix A.6.
-We notice that the multilingual OpenFlamingo trained additionally on mOSCAR gets better results than the model trained on captioning data only while having seen fewer image-text pairs during training.
-More importantly, when increasing the number of few-shot examples from 0 to 16, it sees gains of on average +8.19 points on VQA benchmarks and +16.07 CideR points on captioning benchmarks.
-In contrast, the model trained on text-image pairs only sees gains of +2.82 and +9.08 points respectively.
-In cross-modal machine translation, the model additionally trained on interleaved data is again far better than the one trained on just captioning data, which is not able to translate the Multi30k benchmark at all.
-[^15] Most of the time, the model is not able to follow the prompt and only outputs the end of sequence token.
-Moreover, mOSCAR helps the model to learn to zero-shot disambiguate translations as shown by the improved average score on CoMMuTE (63.75) compared to the model trained on captions only (61.36).
-Multilingual Open Flamingo trained on mOSCAR & text-image pairs is also better than OpenFlamingo 3B MPT evaluated on translate test benchmarks.
-However, we obtain the best results (except for MaXM) by evaluating our multilingual Open Flamingo on the translate-test benchmarks since the underlying language model (Gemma-2B) is far better in English than other languages.
-We also notice that all models struggle with reasoning classification tasks (MaRVL, XVNLI) where they obtain scores close to random guessing.
-xFlickR&CO XM3600 xGQA MaXM MaRVL XVNLI Multi30K CoMMuTE
-#shots
-Multi. OF 0 19.07 8.73 25.08 19.64 49.77 33.01 22.70 63.75
-4 34.32 20.59 31.90 23.90 49.67 36.07 22.79 63.65
-full 8 36.77 22.15 33.9 24.41 49.72 37.16 23.21 63.00
-16 37.63 22.24 35.71 25.38 49.73 35.36 23.48 62.77
-Multi. OF 0 9.57 4.21 8.62 4.01 49.88 33.76 00.00 61.36
-4 13.20 9.26 13.45 4.15 49.54 32.04 00.00 61.13
-cap. only 8 18.00 10.35 12.82 4.88 49.65 33.71 00.01 60.90
-16 19.87 12.07 13.37 4.89 49.79 32.70 00.74 60.25
-Table 4: Results averaged over all languages.
 xGQA MaXM MaRVL XVNLI
@@
 4 35.55 9.89 48.99 36.10
-full 8 36.78 10.12 50.54 39.69
+mOSCAR + cap. 8 36.78 10.12 50.54 39.69
 16 37.75 11.49 49.57 37.97
-Table 5: Translate-test results averaged over languages.
-Figure 4: Score differences averaged over benchmarks and languages between the model trained on mOSCAR + text-image pairs and the model trained only on text-image pairs.
+Table 8: Translate-test results averaged over languages where all benchmarks were translated from local languages into English using Google Translate API.
+Multi.
+OF mOSCAR + cap.
+refers to Multilingual Open Flamingo trained on mOSCAR and text-image pairs while OF-3B MPT refers to Open Flamingo (Awadalla et al., 2023) based on MPT (Team, 2023) and trained on mmc4 (Zhu et al., 2023) and English text-image pairs.
+Figure 4: Score differences averaged over benchmarks and languages between the model trained on mOSCAR + text-image pairs and the model trained only on text-image pairs. Bold is best result.
+5.2 Results
+Tables 5 and 8 show the average results across all languages.
+Full results are available in Appendix A.7.
+We notice that the multilingual OpenFlamingo trained additionally on mOSCAR gets better results than the model trained on captioning data only while having seen fewer image-text pairs during training.
+More importantly, when increasing the number of few-shot examples from 0 to 16, it sees gains of on average +6.71 points on VQA benchmarks and +19.39 CideR points on captioning benchmarks.
+In contrast, the model trained on text-image pairs only sees gains of +2.82 and +9.08 points respectively.
+In cross-modal machine translation, the model additionally trained on interleaved data is again much better than the one trained on just captioning data, which is not able to translate the Multi30k benchmark at all.
+[^20] Most of the time, the model is not able to follow the prompt and only outputs the end of sequence token.
+Moreover, mOSCAR helps the model to learn to zero-shot disambiguate translations as shown by the improved average score on CoMMuTE (63.54) compared to the model trained on captions only (61.36).
+Multilingual Open Flamingo trained on mOSCAR & text-image pairs is also better than OpenFlamingo 3B MPT evaluated on translate test benchmarks [^21] This means benchmarks were translated from local languages to English, using Google Translate API.
+However, we obtain the best results (except for MaXM) by evaluating our multilingual Open Flamingo on the translate-test benchmarks since the underlying language model (Gemma-2B) is far better in English than other languages.
+We also notice that all models struggle with reasoning classification tasks (MaRVL, XVNLI) where they obtain scores close to random guessing.
+Table 6 additionally shows that Multilingual Open Flamingo trained on mOSCAR obtains much better results than the same model trained on WIT for equivalent training data seen during training [^22] We select the checkpoint of multilingual Open Flamingo trained on 35M documents and 70M captions to have fair comparison.
+(except for Multi30K benchmark) which means mOSCAR is better suited than WIT for training multilingual mLLMs.
+Eventually, Table 7 shows that we don’t face a drop in performances in English performances when training the model on 43 languages (multilingual Open Flamingo) in comparison to training it on the English subset of mOSCAR and English text-image pairs.
+Additional comparison results with InternVL2 (Chen et al., 2024), Llava-NeXT (Li et al., 2024), PaliGemma (Beyer* et al., 2024) and Idefics2 (Laurençon et al., 2024) can be found in Appendix A.8.
 We additionally compare results at different training steps, defined by the number of images seen during training.
@@
 Particularly, the gap is wider for few-shot learning.
-6 Conclusion, Limitations and Societal Impacts
-We introduce mOSCAR, a large-scale multilingual and multimodal dataset covering 163 languages and composed of 315M documents, 214B tokens and 1.2B images.
-We show that mOSCAR is of good quality, diverse and can be used to train a multilingual and multimodal LLM.
-We ensure that mOSCAR is as safe as possible by applying a series of filtering steps to remove NSFW content.
-We however did not conduct any toxicity analysis or evaluate its biases as this is challenging in a multilingual setting.
-As it is crawled from the internet, it is possible that mOSCAR reflects biases widespread on it.
-Nevertheless, by its multilingual nature, mOSCAR is a step towards the inclusion of more languages, cultures and people in accessing mLLMs.
+6 Conclusion
+We introduced mOSCAR, a large-scale multilingual and multimodal dataset covering 163 languages and composed of 303M documents, 200B tokens and 1.15B images.
+We showed that mOSCAR is of good quality, diverse and could be used to train a multilingual and multimodal LLM.
+We also proved that training on mOSCAR led to the emergence of in-context learning capabilities in several languages.
+We eventually ensured that mOSCAR is as safe as possible by applying a series of filtering steps to remove NSFW and toxic content.
+Limitations
+We did not conduct any analysis of the biases of mOSCAR as this is challenging in a multilingual setting.
+As it is crawled from the Internet, it is indeed possible that mOSCAR reflects biases widespread on it.
+Training a model on mOSCAR must therefore be combined with additional alignment training steps to mitigate potential biases towards groups of people.
+Nevertheless, by its multilingual nature, mOSCAR is a step towards the inclusion of more languages, cultures, and people in accessing mLLMs.
 Acknowledgements
@@
 We deeply thanks the Jean-Zay support team.
-We also thanks Filip Šedivý for insightful discussions regarding the removal of CSAM, Thorn for having provided access to their CSAM detector, Zeeshan Khan for discussions regarding the training of the models and Victoria Le Fourner for having manually checked subsamples of NSFW images.
+We also thank Filip Šedivý for insightful discussions regarding the removal of CSAM, Thorn for having provided access to their CSAM detector, Zeeshan Khan for discussions regarding the training of the models and Victoria Le Fourner for having manually checked subsamples of NSFW images.
 Appendix A Appendix
 A.1 mOSCAR languages & statistics
+Table LABEL:langs_stats shows the number of documents, the number of images, and the number of NLLB tokens for each language of mOSCAR.
+A.2 Topic modeling
+We analyzed the topic found in different language subsets of mOSCAR with BERTopic (Grootendorst, 2022).
+We first preprocessed $\sim$30K randomly sampled documents in tokens using stanza tokenizers (Qi et al., 2020), we then represent each document with their document embedding obtained with a multilingual sentence-bert232323sentence-transformers/paraphrase-multilingual-MiniLM-L12-v2 (Reimers and Gurevych, 2019) before running BERTopic to cluster the documents.
+Tables 10, 11, 12, 13 and 14 show some of the most salient English-translated strings representing each cluster and their weight (in percentage) in the subsets of mOSCAR.
+While many clusters are shared between subsets, we can observe that some of them are culturally-specific (i.e.
+they are related to the culture of the language subset and only appears on this subset).
+For instance, Table 14 shows topics of the Hindi subset and we observe a cluster related to Hindu gods described by the words ‘ganesha’, ‘shiva’ or ‘shani dev’.
+We also observe a cluster related to festivals hosted in India, it is described by the words ‘festival’, ‘dussehra’ or ‘janmashtami’.
+We can also observe that among clusters shared between several language subsets, some contain knowledge culturally-specific as exemplified by the cluster related to jewelry in the Hindi subset which contain the word ‘mangalsutra’ which is a jewel worn by men during weddings in India.
+Eventually, we can observe that the English subset contains a strong cluster related to the Christian religion.
+While a similar cluster appears in the Tosk Albanian subset, it is related to Islam.
+This shows the importance of a multilingual corpora to have data representing more cultures around the world.
+A.3 Examples of documents
+Figures 5, 7, 7, 8, 9 and 10 provide examples of documents found in mOSCAR for different languages.
+We can observe the diversity in terms of content of the documents.
+Figure 5 is a French document describing a specific shot of golf illustrated by images which provide very detailed description of images.
+Figure 7 shows a Russian document describing an Instagram account with pictures of food with animal heads which is reminiscent of a famous problem of computer vision consisting in classifying images between chihuahuas and muffins.
 Languages Statistics
 Lang. name Code Family Script #documents #images #tokens
-Acehnese ace_Latn Austronesian Latin 7,803 32,461 2,889,134
-Mesopotamian Arabic acm_Arab Afro-Asiatic Arabic 2,274 10,620 1,047,748
-Tunisian Arabic aeb_Arab Afro-Asiatic Arabic 7,640 41,570 2,715,187
-Afrikaans afr_Latn Indo-European Latin 54,895 247,774 39,956,585
-South Levantine Arabic ajp_Arab Afro-Asiatic Arabic 12,098 87,837 5,167,813
-Tosk Albanian als_Latn Indo-European Latin 861,678 2,569,164 452,737,251
-Amharic amh_Ethi Afro-Asiatic Ge‘ez 39,588 152,646 35,089,019
-North Levantine Arabic apc_Arab Afro-Asiatic Arabic 19,904 128,966 9,560,701
-Modern Standard Arabic arb_Arab Afro-Asiatic Arabic 3,936,851 15,126,931 3,401,919,964
-Najdi Arabic ars_Arab Afro-Asiatic Arabic 60,229 296,741 43,610,873
-Moroccan Arabic ary_Arab Afro-Asiatic Arabic 142,386 698,051 204,723,454
-Egyptian Arabic arz_Arab Afro-Asiatic Arabic 835,529 4,054,632 653,626,387
-Assamese asm_Beng Indo-European Bengali 3,948 9,210 640,390
-Asturian ast_Latn Indo-European Latin 165,745 962,723 37,547,944
-Awadhi awa_Deva Indo-European Devanagari 29,324 107,483 4,961,635
-Central Aymara ayr_Latn Aymaran Latin 27,384 151,889 5,148,970
-South Azerbaijani azb_Arab Turkic Arabic 8,274 38,233 5,256,693
-North Azerbaijani azj_Latn Turkic Latin 516,021 1,808,060 257,825,849
-Bashkir bak_Cyrl Turkic Cyrillic 4,532 17,174 3,038,766
-Bambara bam_Latn Manding Latin 7,674 39,190 1,243,332
-Balinese ban_Latn Austronesian Latin 1,886 11,266 542,015
-Belarusian bel_Cyrl Indo-European Cyrillic 63,309 287,539 72,976,520
-Bemba bem_Latn Atlantic–Congo Latin 1,096 7,479 1,340,471
-Bengali ben_Beng Indo-European Bengali 270,406 947,035 35,858,814
-Bhojpuri bho_Deva Indo-European Devanagari 6,366 28,131 875,463
-Banjar bjn_Latn Austronesian Latin 5,427 27,803 1,898,526
-Bosnian bos_Latn Indo-European Latin 1,960,599 7,633,049 1,255,000,505
-Buginese bug_Latn Austronesian Latin 3,312 18,648 588,678
-Bulgarian bul_Cyrl Indo-European Cyrillic 2,591,998 11,670,028 1,760,971,620
-Catalan cat_Latn Indo-European Latin 1,153,864 4,736,634 606,447,390
-Cebuano ceb_Latn Austronesian Latin 16,990 91,234 10,748,818
-Czech ces_Latn Indo-European Latin 3,918,837 13,291,309 2,823,172,996
-Central Kurdish ckb_Arab Indo-European Arabic 36,725 136,566 22,322,689
-Crimean Tatar crh_Latn Turkic Latin 6,376 24,124 1,742,727
-Welsh cym_Latn Indo-European Latin 40,408 165,897 27,748,345
-Danish dan_Latn Indo-European Latin 2,076,298 9,559,600 1,238,277,499
-German deu_Latn Indo-European Latin 20,662,696 87,976,200 8,544,986,218
-Southwestern Dinka dik_Latn Nilo-Saharan Latin 1,712 6,635 1,319,943
-Greek ell_Grek Indo-European Greek 4,916,081 15,209,058 2,923,201,041
-English eng_Latn Indo-European Latin 52,215,013 207,904,315 33,570,108,782
-Esperanto epo_Latn Artificial Latin 25,157 124,996 28,586,195
-Estonian est_Latn Uralic Latin 1,040,368 5,217,366 619,215,048
-Basque eus_Latn Isolate Latin 849,043 3,445,539 277,145,498
-Faroese fao_Latn Indo-European Latin 15,411 60,340 6,691,327
-Fijian fij_Latn Austronesian Latin 1,528 8,776 487,388
-Finnish fin_Latn Uralic Latin 2,396,033 10,365,333 1,781,044,864
-French fra_Latn Indo-European Latin 20,305,739 78,179,601 14,362,579,829
-Friulian fur_Latn Indo-European Latin 37,290 256,456 5,949,600
-Nigerian Fulfulde fuv_Latn Atlantic-Congo Latin 1,568 7,124 401,852
-West Central Oromo gaz_Latn Afro-Asiatic Latin 4,058 11,763 1,786,093
-Scottish Gaelic gla_Latn Indo-European Latin 29,710 153,249 14,605,090
-Irish gle_Latn Indo-European Latin 68,858 315,132 47,438,400
-Galician glg_Latn Indo-European Latin 518,973 2,381,475 217,063,180
-Guarani grn_Latn Tupian Latin 490,945 2,416,633 89,921,114
-Gujarati guj_Gujr Indo-European Gujarati 23,062 91,320 3,324,866
-Haitian Creole hat_Latn Indo-European Latin 257,745 1,570,699 62,847,106
-Hausa hau_Latn Afro-Asiatic Latin 25,364 104,934 13,089,932
-Hebrew heb_Hebr Afro-Asiatic Hebrew 1,109,591 4,766,483 893,327,320
-Hindi hin_Deva Indo-European Devanagari 579,430 1,830,667 122,558,353
-Chhattisgarhi hne_Deva Indo-European Devanagari 1,581 7,263 273,174
-Croatian hrv_Latn Indo-European Latin 1,719,617 8,425,510 1,010,674,096
-Hungarian hun_Latn Uralic Latin 3,534,506 15,390,083 2,831,715,050
-Armenian hye_Armn Indo-European Armenian 339,962 1,141,885 205,635,952
-Igbo ibo_Latn Atlantic-Congo Latin 11,529 68,049 8,701,070
-Ilocano ilo_Latn Austronesian Latin 78,872 523,195 8,116,113
-Indonesian ind_Latn Austronesian Latin 7,016,291 17,324,777 3,981,843,468
-Icelandic isl_Latn Indo-European Latin 244,676 1,027,465 137,015,973
-Italian ita_Latn Indo-European Latin 12,937,153 47,476,971 8,311,790,842
-Javanese jav_Latn Austronesian Latin 24,785 135,583 16,908,805
-Japanese jpn_Jpan Japonic Kanji 14,415,292 23,893,768 8,923,348,944
-Kabyle kab_Latn Afro-Asiatic Latin 18,508 106,730 4,079,553
-Kannada kan_Knda Dravidian Kannada 12,978 42,621 1,442,776
-Kashmiri kas_Arab Indo-European Arabic 3,109 11,408 5,731,910
-Georgian kat_Geor Kartvelian Georgian 354,436 1,304,281 275,223,026
-Kazakh kaz_Cyrl Turkic Cyrillic 252,242 732,648 140,049,214
-Halh Mongolian khk_Cyrl Mongolic Cyrillic 124,412 508,217 84,535,241
-Khmer khm_Khmr Austroasiatic Kher 24,495 122,243 3,043,925
-Kinyarwanda kin_Latn Atlantic-Congo Latin 30,401 172,201 12,049,616
-Kyrgyz kir_Cyrl Uralic Cyrillic 53,010 199,713 34,404,281
-Northern Kurdish kmr_Latn Indo-European Latin 39,262 164,666 23,834,960
-Korean kor_Hang Koreanic Hanja 2,614,089 13,563,283 2,006,080,705
-Lao lao_Laoo Kra-Dai Lao 50,611 208,768 31,029,380
-Ligurian lij_Latn Indo-European Latin 8,751 56,266 2,958,179
-Limburgish lim_Latn Indo-European Latin 189,547 1,076,047 42,534,327
-Lingala lin_Latn Atlantic-Congo Latin 24,614 152,132 4,053,459
-Lithuanian lit_Latn Indo-European Latin 1,688,811 8,869,443 1,161,476,040
-Lombard lmo_Latn Indo-European Latin 30,506 151,855 9,058,614
-Latgalian ltg_Latn Indo-European Latin 11,948 61,624 4,148,492
-Luxembourgish ltz_Latn Indo-European Latin 44,987 246,346 16,676,872
-Ganda lug_Latn Afro-Asiatic Latin 1,878 7,215 789,917
-Mizo lus_Latn Sino-Tibetan Latin 7,880 26,817 4,978,472
-Standard Latvian lvs_Latn Indo-European Latin 896,243 4,141,648 587,653,855
-Magahi mag_Deva Indo-European Devanagari 1,097 3,847 205,763
-Malayalam mal_Mlym Dravidian Malayalam 14,140 52,679 1,689,010
-Marathi mar_Deva Indo-European Devanagari 50,391 163,868 6,689,250
-Minangkabau min_Latn Austronesian Latin 9,341 35,309 1,256,931
-Macedonian mkd_Cyrl Indo-European Cyrillic 542,250 1,853,070 307,232,151
-Maltese mlt_Latn Afro-Asiatic Latin 120,888 709,242 36,097,957
-Maori mri_Latn Austronesian Latin 24,322 130,137 24,957,914
-Burmese mya_Mymr Sino-Tibetan Mon 8,144 44,188 539,527
-Dutch nld_Latn Indo-European Latin 17,096,727 65,606,013 9,670,041,731
-Norwegian Nynorsk nno_Latn Indo-European Latin 199,355 1,012,313 67,799,774
-Norwegian Bokmål nob_Latn Indo-European Latin 2,229,702 9,698,128 1,294,178,095
-Nepali npi_Deva Indo-European Devanagari 31,239 127,193 3,138,539
-Nyanja nya_Latn Atlantic-Congo Latin 12,047 67,192 8,596,769
-Occitan oci_Latn Indo-European Latin 164,852 671,881 59,309,549
-Odia ory_Orya Indo-European Odia 4,319 15,574 378,635
-Pangasinan pag_Latn Austronesian Latin 4,214 32,287 546,071
-Eastern Panjabi pan_Guru Indo-European Gurmukhi 11,497 46,168 1,887,991
-Papiamento pap_Latn Indo-European Latin 55,224 363,015 10,002,655
-Southern Pasto pbt_Arab Indo-European Arabic 32,604 110,807 29,170,322
-Western Persian pes_Arab Indo-European Arabic 7,048,946 25,200,571 6,210,479,015
-Plateau Malgasy plt_Latn Austronesian Latin 32,521 120,673 29,263,848
-Polish pol_Latn Indo-European Latin 14,549,605 60,639,244 11,104,144,109
-Portuguese por_Latn Indo-European Latin 8,145,664 26,530,423 4,760,063,083
-Dari prs_Arab Indo-European Arabic 515,041 2,589,859 517,053,967
-Ayacucho Quechua quy_Latn Quechuan Latin 1,578 11,817 362,690
-Romanian ron_Latn Indo-European Latin 5,180,171 17,964,048 3,548,291,261
-Rundi run_Latn Atlantic-Congo Latin 20,001 67,096 8,686,054
-Russian rus_Cyrl Indo-European Cyrillic 15,913,845 69,542,828 18,909,213,208
-Sango sag_Latn Atlantic-Congo Latin 2,124 13,556 454,455
-Sicilian scn_Latn Indo-European Latin 73,199 424,362 27,110,743
-Sinhala sin_Sinh Indo-European Sinhalese 58,767 221,183 14,270,972
-Slovak slk_Latn Indo-European Latin 3,008,599 15,067,234 1,963,804,563
-Slovenian slv_Latn Indo-European Latin 1,472,025 7,210,285 935,834,754
-Samoan smo_Latn Austronesian Latin 12,346 71,359 14,954,824
-Shona sna_Latn Atlantic-Congo Latin 12,698 68,782 6,112,600
-Sindhi snd_Arab Indo-European Arabic 21,095 74,289 17,647,825
-Somali som_Latn Afro-Asiatic Latin 77,343 301,429 34,554,975
-Southern Sotho sot_Latn Atlantic-Congo Latin 7,718 43,146 6,156,450
-Spanish spa_Latn Indo-European Latin 22,713,366 78,361,087 14,616,773,475
-Sardinian srd_Latn Indo-European Latin 675,539 4,059,493 106,159,957
-Serbian srp_Cyrl Indo-European Cyrillic 604,557 2,286,171 401,223,741
-Sundanese sun_Latn Austronesian Latin 44,310 236,025 13,627,832
-Swedish swe_Latn Indo-European Latin 3,302,730 10,860,518 1,779,284,152
-Swahili swh_Latn Atlantic-Congo Latin 137,134 593,418 59,454,896
-Silesian szl_Latn Indo-European Latin 23,535 132,459 5,996,972
-Tamil tam_Taml Dravidian Tamil 36,196 167,669 4,834,946
-Tatar tat_Cyrl Turkic Cyrillic 37,188 143,842 22,831,350
-Telugu tel_Telu Dravidian Telugu 22,974 81,033 2,273,772
-Tajik tgk_Cyrl Turkic Cyrillic 125,236 417,591 90,503,778
-Tagalog tgl_Latn Austronesian Latin 151,437 673,814 97,708,639
-Thai tha_Thai Kra-Dai Thai 2,983,837 11,621,786 2,839,211,104
-Tigrinya tir_Ethi Afro-Asiatic Ge‘ez 2,657 8,707 1,725,422
-Tok Pisin tpi_Latn Indo-European Latin 5,063 35,169 460,853
-Turkmen tuk_Latn Turkic Latin 13,024 57,354 9,766,999
-Turkish tur_Latn Turkic Latin 4,478,700 12,401,091 2,394,669,068
-Twi twi_Latn Atlantic-Congo Latin 3,305 13,634 495,220
-Uyghur uig_Arab Turkic Arabic 10,713 41,709 6,785,318
-Ukrainian ukr_Cyrl Indo-European Cyrillic 2,721,424 10,929,796 1,928,351,595
-Urdu urd_Arab Indo-European Arabic 407,098 1,239,125 242,007,283
-Northern Uzbek uzn_Latn Turkic Latin 156,632 798,155 89,022,562
-Venetian vec_Latn Indo-European Latin 330,611 1,830,777 71,077,531
-Vietnamese vie_Latn Viet-Muong Latin 12,621,521 47,411,488 11,616,191,199
-Wolof wol_Latn Atlantic-Congo Latin 4,658 20,380 1,596,432
-Xhosa xho_Latn Atlantic-Congo Latin 25,950 142,387 15,809,823
-Eastern Yiddish ydd_Hebr Indo-European Hebrew 12,486 57,510 17,369,727
-Yoruba yor_Latn Atlantic-Congo Latin 56,700 286,933 32,614,558
-Yue Chinese yue_Hant Sino-Tibetan Hant 33,671 203,513 24,172,441
-Chinese (Simplified) zho_Hans Sino-Tibetan Hanzi 9,861,262 36,152,754 8,078,842,701
-Chinese (Traditional) zho_Hant Sino-Tibetan Hant 3,967,966 16,307,258 2,962,854,441
-Standard Malay zsm_Latn Austronesian Latin 1,179,744 5,488,632 432,667,199
-Zulu zul_Latn Atlantic-Congo Latin 30,717 156,639 11,345,288
+Acehnese ace_Latn Austronesian Latin 2,159 9,026 1,395,381
+Mesopotamian Arabic acm_Arab Afro-Asiatic Arabic 1,282 5,621 704,549
+Tunisian Arabic aeb_Arab Afro-Asiatic Arabic 5,933 34,270 2,308,455
+Afrikaans afr_Latn Indo-European Latin 50,061 211,876 38,761,504
+South Levantine Arabic ajp_Arab Afro-Asiatic Arabic 8,603 69,051 3,869,688
+Tosk Albanian als_Latn Indo-European Latin 856,144 2,543,758 441,244,377
+Amharic amh_Ethi Afro-Asiatic Ge‘ez 39,031 149,739 33,768,732
+North Levantine Arabic apc_Arab Afro-Asiatic Arabic 16,198 110,792 8,268,237
+Modern Standard Arabic arb_Arab Afro-Asiatic Arabic 3,794,792 14,757,353 3,346,786,610
+Najdi Arabic ars_Arab Afro-Asiatic Arabic 52,102 261,275 39,066,487
+Moroccan Arabic ary_Arab Afro-Asiatic Arabic 117,957 584,301 188,462,338
+Egyptian Arabic arz_Arab Afro-Asiatic Arabic 761,113 3,785,164 635,018,784
+Assamese asm_Beng Indo-European Bengali 2,947 7,228 543,676
+Asturian ast_Latn Indo-European Latin 87,649 533,723 25,499,269
+Awadhi awa_Deva Indo-European Devanagari 8,179 29,142 2,293,620
+Central Aymara ayr_Latn Aymaran Latin 10,112 57,294 2,343,403
+South Azerbaijani azb_Arab Turkic Arabic 3,411 14,825 3,143,946
+North Azerbaijani azj_Latn Turkic Latin 511,832 1,796,046 256,160,442
+Bashkir bak_Cyrl Turkic Cyrillic 3,287 12,031 2,600,135
+Bambara bam_Latn Manding Latin 3,011 17,666 446,961
+Balinese ban_Latn Austronesian Latin 787 4,894 392,978
+Belarusian bel_Cyrl Indo-European Cyrillic 60,443 276,672 71,854,171
+Bemba bem_Latn Atlantic–Congo Latin 582 3,018 1,021,026
+Bengali ben_Beng Indo-European Bengali 204,475 758,222 30,400,395
+Bhojpuri bho_Deva Indo-European Devanagari 4,190 18,339 715,786
+Banjar bjn_Latn Austronesian Latin 1,764 9,017 1,093,443
+Bosnian bos_Latn Indo-European Latin 635,750 2,642,491 423,073,661
+Buginese bug_Latn Austronesian Latin 584 2,379 167,459
+Bulgarian bul_Cyrl Indo-European Cyrillic 2,578,191 11,601,214 1,736,106,287
+Catalan cat_Latn Indo-European Latin 1,132,056 4,638,966 598,942,711
+Cebuano ceb_Latn Austronesian Latin 14,924 75,258 10,221,371
+Czech ces_Latn Indo-European Latin 3,736,126 12,683,461 2,767,295,966
+Central Kurdish ckb_Arab Indo-European Arabic 36,413 135,461 21,622,335
+Crimean Tatar crh_Latn Turkic Latin 2,744 10,079 1,173,321
+Welsh cym_Latn Indo-European Latin 38,616 155,591 27,237,252
+Danish dan_Latn Indo-European Latin 2,020,516 9,214,031 1,207,829,704
+German deu_Latn Indo-European Latin 20,265,504 86,393,702 8,315,212,019
+Southwestern Dinka dik_Latn Nilo-Saharan Latin 1,233 4,766 1,098,795
+Greek ell_Grek Indo-European Greek 4,895,433 15,147,284 2,909,427,055
+English eng_Latn Indo-European Latin 51,658,029 205,363,181 32,599,001,993
+Esperanto epo_Latn Artificial Latin 23,619 112,577 26,976,847
+Estonian est_Latn Uralic Latin 1,022,368 5,108,102 589,045,973
+Basque eus_Latn Isolate Latin 682,599 2,914,120 259,930,954
+Faroese fao_Latn Indo-European Latin 14,921 56,934 6,579,921
+Fijian fij_Latn Austronesian Latin 1,039 4,039 416,670
+Finnish fin_Latn Uralic Latin 2,377,155 10,263,171 1,749,904,041
+French fra_Latn Indo-European Latin 19,963,542 76,851,982 13,818,099,493
+Friulian fur_Latn Indo-European Latin 15,823 120,878 2,550,209
+Nigerian Fulfulde fuv_Latn Atlantic-Congo Latin 919 4,281 264,234
+West Central Oromo gaz_Latn Afro-Asiatic Latin 3,399 9,071 1,640,693
+Scottish Gaelic gla_Latn Indo-European Latin 19,638 105,937 13,119,348
+Irish gle_Latn Indo-European Latin 60,303 267,562 45,341,371
+Galician glg_Latn Indo-European Latin 410,489 1,696,763 197,685,077
+Guarani grn_Latn Tupian Latin 207,800 1,038,296 48,610,979
+Gujarati guj_Gujr Indo-European Gujarati 21,916 87,805 3,202,096
+Haitian Creole hat_Latn Indo-European Latin 105,777 667,801 34,261,838
+Hausa hau_Latn Afro-Asiatic Latin 21,850 81,141 11,807,898
+Hebrew heb_Hebr Afro-Asiatic Hebrew 1,098,800 4,708,947 859,238,720
+Hindi hin_Deva Indo-European Devanagari 543,928 1,745,222 118,903,998
+Chhattisgarhi hne_Deva Indo-European Devanagari 832 3,908 205,345
+Croatian hrv_Latn Indo-European Latin 1,689,553 8,315,237 998,928,993
+Hungarian hun_Latn Uralic Latin 3,515,058 15,293,132 2,811,446,583
+Armenian hye_Armn Indo-European Armenian 336,285 1,126,920 199,883,484
+Igbo ibo_Latn Atlantic-Congo Latin 7,089 41,672 3,014,602
+Ilocano ilo_Latn Austronesian Latin 7,076 59,327 832,454
+Indonesian ind_Latn Austronesian Latin 6,644,918 16,237,247 2,895,956,979
+Icelandic isl_Latn Indo-European Latin 239,195 1,003,522 131,308,802
+Italian ita_Latn Indo-European Latin 12,812,932 47,011,085 8,144,757,759
+Javanese jav_Latn Austronesian Latin 18,192 100,952 15,206,708
+Japanese jpn_Jpan Japonic Kanji 14,154,575 23,435,549 8,539,956,266
+Kabyle kab_Latn Afro-Asiatic Latin 6,101 33,923 1,781,992
+Kannada kan_Knda Dravidian Kannada 9,373 33,147 1,206,651
+Kashmiri kas_Arab Indo-European Arabic 1,498 5,284 3,384,394
+Georgian kat_Geor Kartvelian Georgian 353,471 1,300,710 274,042,522
+Kazakh kaz_Cyrl Turkic Cyrillic 248,403 718,126 138,597,176
+Halh Mongolian khk_Cyrl Mongolic Cyrillic 123,789 505,098 83,628,495
+Khmer khm_Khmr Austroasiatic Kher 23,348 116,437 2,915,205
+Kinyarwanda kin_Latn Atlantic-Congo Latin 20,381 108,280 10,268,334
+Kyrgyz kir_Cyrl Uralic Cyrillic 51,221 194,092 33,981,180
+Northern Kurdish kmr_Latn Indo-European Latin 34,593 142,634 21,972,155
+Korean kor_Hang Koreanic Hanja 2,614,038 13,562,957 2,000,344,511
+Lao lao_Laoo Kra-Dai Lao 49,925 205,452 30,098,274
+Ligurian lij_Latn Indo-European Latin 3,581 26,740 1,046,463
+Limburgish lim_Latn Indo-European Latin 70,099 443,903 25,465,590
+Lingala lin_Latn Atlantic-Congo Latin 6,304 41,400 1,580,536
+Lithuanian lit_Latn Indo-European Latin 1,673,790 8,772,570 1,153,604,941
+Lombard lmo_Latn Indo-European Latin 14,053 61,359 6,270,646
+Latgalian ltg_Latn Indo-European Latin 5,174 21,062 2,903,043
+Luxembourgish ltz_Latn Indo-European Latin 27,946 142,470 13,925,521
+Ganda lug_Latn Afro-Asiatic Latin 1,475 4,118 688,308
+Mizo lus_Latn Sino-Tibetan Latin 7,009 22,630 4,106,536
+Standard Latvian lvs_Latn Indo-European Latin 857,757 3,937,940 578,441,751
+Magahi mag_Deva Indo-European Devanagari 290 1,088 94,031
+Malayalam mal_Mlym Dravidian Malayalam 11,203 44,417 1,420,906
+Marathi mar_Deva Indo-European Devanagari 43,720 142,001 6,164,176
+Minangkabau min_Latn Austronesian Latin 1,523 7,300 447,320
+Macedonian mkd_Cyrl Indo-European Cyrillic 539,149 1,841,846 304,592,615
+Maltese mlt_Latn Afro-Asiatic Latin 56,666 327,331 27,114,870
+Maori mri_Latn Austronesian Latin 20,840 114,680 24,524,962
+Burmese mya_Mymr Sino-Tibetan Mon 6,575 36,661 406,016
+Dutch nld_Latn Indo-European Latin 16,890,074 64,609,055 9,493,533,101
+Norwegian Nynorsk nno_Latn Indo-European Latin 138,384 701,972 57,812,652
+Norwegian Bokmål nob_Latn Indo-European Latin 2,192,012 9,534,178 1,267,421,216
+Nepali npi_Deva Indo-European Devanagari 28,042 116,363 2,892,865
+Nyanja nya_Latn Atlantic-Congo Latin 11,749 65,324 8,513,823
+Occitan oci_Latn Indo-European Latin 61,681 323,632 21,029,975
+Odia ory_Orya Indo-European Odia 3,759 14,373 340,695
+Pangasinan pag_Latn Austronesian Latin 1,045 7,770 270,363
+Eastern Panjabi pan_Guru Indo-European Gurmukhi 10,857 44,440 1,821,511
+Papiamento pap_Latn Indo-European Latin 29,564 177,229 7,396,392
+Southern Pashto pbt_Arab Indo-European Arabic 31,854 107,563 27,623,486
+Western Persian pes_Arab Indo-European Arabic 6,995,368 24,998,370 6,061,794,870
+Plateau Malgasy plt_Latn Austronesian Latin 32,119 119,506 28,542,084
+Polish pol_Latn Indo-European Latin 14,492,239 60,362,860 10,994,239,010
+Portuguese por_Latn Indo-European Latin 8,033,406 26,058,040 4,639,089,792
+Dari prs_Arab Indo-European Arabic 421,097 2,101,038 399,037,437
+Ayacucho Quechua quy_Latn Quechuan Latin 1,248 10,038 322,112
+Romanian ron_Latn Indo-European Latin 5,131,444 17,790,793 3,484,865,185
+Rundi run_Latn Atlantic-Congo Latin 17,798 55,060 8,140,230
+Russian rus_Cyrl Indo-European Cyrillic 15,753,144 68,786,134 18,196,141,357
+Sango sag_Latn Atlantic-Congo Latin 724 4,564 181,876
+Sicilian scn_Latn Indo-European Latin 27,388 164,772 17,535,500
+Sinhala sin_Sinh Indo-European Sinhalese 44,963 179,082 11,413,044
+Slovak slk_Latn Indo-European Latin 2,979,681 14,894,160 1,951,406,321
+Slovenian slv_Latn Indo-European Latin 1,456,026 7,106,291 928,101,642
+Samoan smo_Latn Austronesian Latin 11,024 62,358 11,672,900
+Shona sna_Latn Atlantic-Congo Latin 7,400 41,385 5,276,139
+Sindhi snd_Arab Indo-European Arabic 20,615 70,992 16,686,668
+Somali som_Latn Afro-Asiatic Latin 58,151 209,905 31,093,227
+Southern Sotho sot_Latn Atlantic-Congo Latin 7,474 41,714 5,876,842
+Spanish spa_Latn Indo-European Latin 22,218,630 76,372,709 13,882,047,139
+Sardinian srd_Latn Indo-European Latin 336,476 2,220,976 68,281,992
+Serbian srp_Cyrl Indo-European Cyrillic 593,332 2,251,042 394,477,097
+Sundanese sun_Latn Austronesian Latin 16,438 89,379 9,549,957
+Swedish swe_Latn Indo-European Latin 3,231,753 10,558,719 1,748,495,813
+Swahili swh_Latn Atlantic-Congo Latin 96,770 365,792 52,827,863
+Silesian szl_Latn Indo-European Latin 7,846 47,313 3,022,502
+Tamil tam_Taml Dravidian Tamil 30,202 149,837 4,234,345
+Tatar tat_Cyrl Turkic Cyrillic 34,489 133,014 22,255,423
+Telugu tel_Telu Dravidian Telugu 16,107 54,100 1,633,579
+Tajik tgk_Cyrl Turkic Cyrillic 119,383 395,470 87,519,228
+Tagalog tgl_Latn Austronesian Latin 140,922 628,210 95,285,900
+Thai tha_Thai Kra-Dai Thai 1,799,735 6,603,060 807,374,946
+Tigrinya tir_Ethi Afro-Asiatic Ge‘ez 2,622 8,601 1,699,272
+Tok Pisin tpi_Latn Indo-European Latin 785 5,888 97,298
+Turkmen tuk_Latn Turkic Latin 12,372 54,002 9,650,172
+Turkish tur_Latn Turkic Latin 4,448,111 12,304,912 2,356,627,784
+Twi twi_Latn Atlantic-Congo Latin 286 2,041 78,227
+Uyghur uig_Arab Turkic Arabic 10,614 41,367 6,602,690
+Ukrainian ukr_Cyrl Indo-European Cyrillic 2,689,369 10,842,572 1,909,330,669
+Urdu urd_Arab Indo-European Arabic 403,245 1,224,175 236,356,788
+Northern Uzbek uzn_Latn Turkic Latin 113,772 581,861 81,808,833
+Venetian vec_Latn Indo-European Latin 122,390 763,029 24,081,966
+Vietnamese vie_Latn Viet-Muong Latin 12,296,989 46,339,341 11,462,111,787
+Wolof wol_Latn Atlantic-Congo Latin 2,152 9,351 367,848
+Xhosa xho_Latn Atlantic-Congo Latin 13,620 80,748 14,566,904
+Eastern Yiddish ydd_Hebr Indo-European Hebrew 12,275 56,421 17,078,751
+Yoruba yor_Latn Atlantic-Congo Latin 10,148 49,474 8,346,193
+Yue Chinese yue_Hant Sino-Tibetan Hant 28,478 172,592 21,579,579
+Chinese (Simplified) zho_Hans Sino-Tibetan Hanzi 8,326,440 29,575,591 5,199,137,981
+Chinese (Traditional) zho_Hant Sino-Tibetan Hant 3,796,336 15,514,804 2,617,463,485
+Standard Malay zsm_Latn Austronesian Latin 864,831 3,651,754 384,708,004
+Zulu zul_Latn Atlantic-Congo Latin 13,089 73,167 9,654,461
 Table 9: Languages & Statistics
-A.2 Heuristics to increase the quality of documents
+Relevant words describing the topics Topic representation (in %)
+recipe, sauce, cheese, recipes, chicken, add, delicious, food, cook, minutes 7.17
+game, games, players, gaming, play, gameplay, pc, playing, player, playstation 2.1
+dog, dogs, cat, pet, cats, pets, puppy, breed, puppies, animal 1.94
+god, jesus, church, christ, faith, lord, bible, him, christian, holy 1.82
+shirt, jacket, jeans, cotton, wear, shirts, fit, men, size, waist 1.68
+estate, property, mortgage, real, home, house, buyers, market, properties, homes 1.6
+art, artist, artists, painting, gallery, paintings, works, arts, his, exhibition 1.59
+card, cards, love, gift, cute, christmas, fun, stamp, halloween, easter 1.57
+covid, 19, vaccine, coronavirus, virus, cases, health, vaccinated, vaccination, vaccines 1.28
+album, band, music, song, songs, rock, release, albums, his 1.27
+data, cloud, server, software, aws, application, business, management, cluster, configuration 1.21
+shipping, item, delivery, ankle, brace, items, order, days 1.14
+diamond, jewelry, ring, carat, gold, rings, earrings, silver, necklace 1.13
+books, book, read, she, reading, author, novel, story 1.13
+dress, fashion, dresses, wear, style, skirt, wedding, outfit, love 1.12
+bedroom, apartments, room, property, apartment, kitchen, floor, home, spacious 1.03
+trump, president, election, biden, republican, democrats, senate, republicans, court 1.0
+bitcoin, crypto, cryptocurrency, blockchain, ethereum, tokens, token, price, cryptocurrencies, nft 0.94
+life, tarot, yourself, love, feel, myself, mind 0.93
+sound, hearing, audio, speakers, noise, headphones, bluetooth, speaker, microphone, wireless 0.9
+trail, lake, park, mountain, hike, trails, hiking, canyon, river, yosemite 0.85
+cookies, website, cookie, settings, resume, preferences, disable, browser, user, information 0.85
+police, said, arrested, man, officers, crime, county 0.84
+skin, acne, skincare, sunscreen, face, treatment, facial, pores, oil, hyamax 0.83
+casino, gambling, games, slot, casinos, online, bonus, poker, slots, players 0.82
+lighting, light, lights, led, lamp, bulb, lamps, bulbs, wall, fixtures 0.76
+league, club, chelsea, season, his, football, cup, players, player 0.76
+wedding, weddings, venue, couples, dj, guests, day, reception, ceremony, bride 0.76
+dental, teeth, tooth, dentist, smile, dentistry, whitening, oral, implant, dentists 0.72
+cbd, cannabis, marijuana, hemp, thc, gummies, cannabinoids, medical, oil, recreational 0.69
+shoes, shoe, boots, socks, nike, sneaker, heel, foot, running, boxing 0.68
+stocks, inflation, stock, market, investment, gold, investors, fund, trading, funds 0.67
+car, cars, engine, rear, porsche, toyota, mercedes, electric, suv 0.65
+fitness, workout, exercise, gym, trainer, exercises, training, strength, body, workouts 0.63
+divorce, attorney, law, lawyer, legal, court, lawyers, attorneys, injury, case 0.59
+steel, nailer, arm, drill, machine, wrench, tool, saw, palm 0.59
+security, cyber, ransomware, cybersecurity, attacks, malware, phishing, attack, data, hackers 0.59
+furniture, chair, chairs, sofa, table, design, style, dining, comfort, leather 0.5
+her, she, actress, star, kristina, instagram, lorena, grikaite, kardashian 0.49
+testosterone, vitamin, supplements, weight, blood, body, supplement, levels, calcium, muscle 0.49
+bag, bags, backpack, stethoscope, pockets, strap, tote, zipper, purse, leather 0.48
+energy, carbon, wind, renewable, emissions, gas, coal, power, electricity, climate 0.47
+ball, tennis, player, sports, sport, players, backhand, coach, athletes, drills 0.47
+phone, samsung, smartphone, camera, galaxy, oneplus, realme, battery, schematic, android 0.47
+watches, replica, watch, rolex, dial, clock, patek, swiss, fake, wholesale 0.47
+fundraising, volunteer, donors, volunteers, donor, charity, charities, community, volunteering, imdsa 0.46
+parking, traffic, pedestrian, rail, transport, transportation, street, mobility, public, city 0.45
+race, f1, racing, lap, formula, drivers, season, car, championship, driver 0.44
+pain, chiropractic, joint, ankle, spinal, muscles, arthritis, spine, symptoms, joints 0.41
+wine, wines, winery, bordeaux, vineyard, tasting, sauvignon, grapes, vineyards, palate 0.41
+Table 10: BERTopic modeling on the English subset of mOSCAR.
+Relevant words describing the topics Topic representation (in %)
+Kosovo, Serbia, European, taken, Kurti, Serbian, Serbia 4.65
+Ukraine, Russian, Russia, Putin, war, Ukrainian, military 4.44
+Covid, 19, vaccine, cases, health, patients, coronavirus 3.72
+Luizi, Kiara, Big, Brother, VIP, Olta, stuck 3.17
+apartment, rent, floor, m2, lease, housing, located, area, sale 2.51
+letter, language, book, literary, poetry, book, Albanian, writer 1.81
+add, spoon, eggs, recipe, sugar, pour, oven, oil, flour 1.46
+photo, she, Instagram, hers, bikini, Tarja, model, follows, dress 1.23
+Allah, greeting, Quran, prophet, interpretation, exegesis, Islam, prophets 1.19
+court, property, court, trials, KPK, reevaluation, appeal, judges, declaration, prosecution 1.16
+horoscope, sign, signs, you, stars, zodiac 0.99
+music, festival, stage, KultPlus, artist, culture, musical 0.98
+prison, court, criminal, imprisonment, pretrial detention, measure, sentenced, crime 0.85
+Inter, Inter Milan, Milan, AC Milan, Gazzetta, Pioli 0.82
+Rama, Edi, prime minister, opposition, PD, McGonigal, Basha, justice, PS 0.82
+Trump, Biden, Donald, president, Republican, president’s, Joe, American, White, president 0.81
+accident, occurred, consequence, road, car, injured, type, crashed 0.78
+Berisha, PD, Sali, Berisha, Rama, opposition, elections, Rama, Edi, party’s 0.74
+exhibition, art, museum, exhibition, culture, artist, KultPlus 0.72
+knife, gold, weapon, police, killed, incident, fired, event, victim 0.7
+water, river, residents, floods, rain, drinking 0.7
+women, gender, violence, LGBT, friend, violence, family 0.68
+migrants, asylum, refugees, permits, Albanian, passport, exchange, asylum seeker, 000, passport 0.63
+students, UET, academic, university, university’s, Erasmus, universities, faculty, study, UBT 0.63
+tourists, tourism, tourist, travel, foreign, destination, Albanian, visitors 0.60
+doctors, patients, health, clinic, hour, sore, clinic, Dr 0.57
+firefighters, fire, flame, fires, extinguishing, hotspots 0.56
+Ronaldo, Cristiano, CR7, Portuguese, United, Manchester, Messi 0.56
+music, musical, Adriano, Celentano, musical, Tari, top, titled 0.56
+weather, temperatures, forecast, cloudy, rain, snow, degrees 0.56
+Mercedes, car, Audi, sale, BMW, Aston, vehicles, Benz, sold 0.52
+email, site, comment, browser, save, address, fields, marked, required 0.5
+archaeologists, archaeological, Roman, ancient, city, archaeological, restored, Durres, century 0.49
+NASA, Earth, space, science, planet, boundary, solar, Apollo, planets 0.47
+Albanian, November, flag, nations, independence, year, day, Vasfije 0.46
+oil, liter, price, fuel, lek prices, gasoline, petrol, barrel 0.45
+PD, Democratic, party’s, party, PS, elections, vote, party, Socialist, Democrat 0.44
+city, located, park, rock, water, natural 0.44
+Turkey, earthquake, Syria, magnitude, collapsed 0.43
+airline, plane, passengers, flight 0.41
+road, buses, urban, bike, works, city’s, urban, municipality, Veliaj, transportation 0.4
+iPhone, Apple, Pro, Max, iOS, iPad, 15, apps, 14, dollars 0.4
+north, North Macedonia, Bulgarian, Bulgaria, Macedonian, Macedonia, Macedonians 0.4
+Table 11: BERTopic modeling on the Tosk Albanian subset of mOSCAR.
+Relevant words describing the topics Topic representation (in %)
+Match, league, Madrid, Barcelona, match, football, Real 9.86
+Allah, Sheikh, Mohammed, occupation 4.24
+Allah, cleaning, homes, jurisprudential, buying, Kuwait, furniture 2.78
+Game, games, download, casino, Android, play, computer 2.5
+Sarai, occupation, Jerusalem, Palestinian, Gaza, next, Palestine 2.08
+Corona, health, virus, infection, Covid, 19, vaccine 1.98
+Ukraine, Russia, Russian, Putin, Moscow, Kiev, forces, Ukrainian 1.86
+Fridge, LED, electric, watts, stainless steel, product, resistant, corrosion 1.66
+Movie, film, TV series, actress, actor, series, cinematic, ceremony, cinema 1.54
+Hospital, medical, health, medical, care, doctors, diseases, doctor 1.47
+Car, cars, Toyota, Hyundai, electric, Mercedes, Ford, Kia 1.44
+Capital, Daraya, compound, real estate, mall, meters, administrative, units, for sale 1.4
+Data, workers, feasibility, project, work, study, customers, employees 1.25
+Syrian, Syria, ISIS, Aleppo, forces 1.01
+Saudi, national, Saudi Arabia, Kingdom, Arabia, Muhammad 0.98
+Aviation, airport, flight, airplane, aircraft, air, airlines 0.91
+Women, women’s, violence, rights, society, feminist, affected, victims, origin 0.83
+iPhone, iPhone, Apple, iOS, phone, iPad, apps 0.83
+Space, NASA, Mars, moon, planet, planets, scientists, telescope 0.81
+Iran, Tehran, nuclear, agreement, Washington, United States, American 0.79
+Tourism, hotel, island, Maldives, beach, Malaysia, shores 0.71
+Skin, cream, face, skin, oil, care, your skin, treatment, oily 0.65
+Chocolate, spoon, meat, recipes, cake, butter, method, recipe, bowl 0.62
+College, engineering, science, education, university, sciences, department, school 0.6
+Temperature, weather, degrees, forecast, airport, temperature, west, regions 0.6
+Police, suspect, arrest, drugs, crime, general, city, prosecutor 0.59
+Spine, column, knee, blood, body, fever, pain, muscle 0.59
+Istanbul, Turkey, Turkish, apartments, for sale, real estate, nationality, Şişli, property 0.59
+Pregnancy, fetus, womb, birth, monthly, cycle, doctor, progesterone, symptoms 0.57
+Sudan, Ethiopia, Tigray, Ethiopian, Sudanese, Ethiopian, Renaissance, army, Sudanese 0.55
+Germany, Sweden, Europe, asylum, Merkel, immigration, refugee 0.54
+Insects, pest control, ants, pests, spray, pesticides, cockroaches, company, white 0.53
+University, Baniyas, scholarships, students, universities, study 0.51
+Gold, price, carat, today, pound, grams, dealings 0.5
+Medicine, drug, prescription, dosage, side effects, doctor, treatment, illness 0.5
+Rooms, bedrooms, decorations, modern, wallpaper, furniture, sleeping 0.49
+Hotels, activities, trips, close, hotel, tours 0.45
+Samsung, Galaxy, phone, specifications, S23, 5G, phones 0.44
+Fire, accident, firefighters, accident, Greece, outbreak, injured, forest 0.42
+Trump, Donald, Biden, elections, American, former, White House 0.41
+Technician, painter, Kuwait, repair, installation, AC, satellite, changing, provided 0.4
+Hair, hair loss, shampoo, scalp, grow, hair, hair treatment 0.39
+Education, secondary, studies, students, exams, schools, cumulative, average 0.39
+Elections, voting, election, Myanmar, council, parties, electoral 0.39
+Buyer, seller, shipment, equipment, Alibaba, help, charger, suppliers, machine, tank 0.39
+India, ambassador, cooperation, Libya, ministers, Modi 0.39
+Table 12: BERTopic modeling on the Arabic subset of mOSCAR.
+Relevant words describing the topics Topic representation (in %)
+team, match, child, event, olympic, goal, championship, win, champion 10.84
+school, students, education, children, second, student 4.11
+elections, syriza, mitsotakis, government, ND, response, prime minister, parliament, no 2.75
+order, shipment, product, payment, purchase, cash on delivery, receipt, donation, you, shipping fees 2.15
+god, god’s, church, holy, christ, priest, prayer, divine 1.8
+gold, necklace, earrings, ring, bracelet, jewelry, silver, steel 1.79
+book, writer, history, published, read, literature, write, novel 1.64
+game, play, children, two, ball, games 1.36
+ukraine, russia, russian, force, usa, war, russia 1.21
+shoes, t-shirt, adidas, nike, sneakers, men, athletic, running 1.18
+song, music, album, single, music, two, rock 1.16
+skin, face, cream, body, oil, oily, skin care 1.15
+traffic, vehicle, road, lane, road, vehicles, street 1.11
+mattress, sofa, pillow, bed, chair, furniture, armchair, bedroom, color 1.04
+recipe, chicken, recipe, juicy, add, chicken, mix, season 1.02
+christmas, birth, christmas, december, gift, festive, tree, holiday 0.98
+turkey, erdogan, turkish, greece, home, democracy, istanbul 0.93
+moon, planet, nasa, sky, galaxy, earth, telescope, satellite, lunar, space 0.83
+tourism, tourist, greece, destination, trips, hotel 0.81
+mayor, march, meeting, municipal, council, committee, session, members 0.81
+ships, boats, port, cruise, shipping, sea, cruise ship, peiraeus 0.78
+fire, fire brigade, wildfire, fire extinguishers, area, blaze, fire vehicles, firemen 0.78
+beach, island, water, village, white, place 0.77
+art, painting, works, artist, museum, exhibition, artists 0.76
+car, bmw, model, new, nissan, audi, mobile, video 0.75
+vaccine, vaccination, covid, doses, health, coronavirus, disease 0.74
+led, lighting, lamp, simple, lights, white, lighting, hanging, ceiling, lamps 0.72
+you, yourself, feeling, life, thoughts, live, feel 0.68
+cotton, cotton, fabric, design, cotton 0.67
+purse, wallet, leather, case, bag, items 0.65
+exercise, gym, workout, practice, body, strength, pilates, muscle 0.64
+weather, rainy, temperature, clouds, cold, storms, celsius, wind 0.63
+company, product, sunlight, quality, service, construction, production, service 0.62
+plants, seeds, plant, grow, flowers, neem, soil, wild 0.62
+europe, euro, greek, eu, policy, government, eurozone, economy 0.61
+wine, wines, vineyard, winery, variety, grapes, alcohol 0.57
+pump, burner, gas, water, temperature, heating, air, water heater 0.55
+headphones, bluetooth, speakers, sound, usb, wireless, audio, sound 0.54
+movie, series, netflix, characters, director, episode, season, naruto 0.54
+apartment, for sale, bedroom, location, new, living, for rent, kitchen 0.49
+restaurant, kitchen, chef, dish, food, menu, restaurant, tavern 0.49
+diet, weight, food, loss, calories, pounds, healthy 0.48
+history, news, breakfast, syriza, greek, government, politics, news 0.47
+video, tv, sony, iptv, dvd, photo, movie, camera 0.47
+dog, dogs, pet, animal, dog, puppies, leash, friend, pets 0.46
+inflation, rates, euro, increase, prices, interest, market, economy 0.42
+Table 13: BERTopic modeling on the Greek subset of mOSCAR.
+Relevant words describing the topics Topic representation (in %)
+loan, bank, paytm, sbi, upi, credit, personal, sunal, atm, account 3.61
+electric, tata, mahindra, hyundai, maruti, bike, suv 3.16
+recipe, make, paneer, chicken, best, ingredients, potato 2.45
+recruitment, notification, availability, 2023, vacancy, age, position, fill 1.39
+whatsapp, sap, fm, gb, chat, gbwhatsapp, message, group, feature 1.28
+movie, khan, film, jawan, trailer, release, series, jaari, cast, collection 1.25
+bitcoin, cryptocurrency, token, crypto, currency, ethereum, coin, blockchain, binance, dogecoin 1.25
+samsung, 5g, oneplus, realme, redmi, nokia, oppo, vivo, galaxy, xiaomi 1.19
+vs, ind, aus, odi, one-day, wicket, match, runs, australia, cup 1.1
+stock, market, trading, intraday, share, ratio, demat, year, buy 0.88
+wish, life, dog, equity, confidence, morning, suprabhat, habit, flirting, self 0.86
+business, ideas, idea, manufacturing, making, small, entrepreneur, start, recycling, people 0.85
+aadhar, aadhaar, uidai, half, card, otp, update, number, pan, bar 0.72
+designs, design, taarak, mehta, alto, jewellery, mangalsutra, galas, blouse, earrings 0.72
+rlalibaba, equipment, mash, ravi, gati, jak, shakash, ens, quot, woodworking 0.69
+iphone, apple, aif, apple, 14, ipad, 15, ios, xs, mac 0.67
+biography, chay, kanika, mann, one, alia, bhatt, tulsidas, family, age 0.63
+reply, cancel, leave, rac, berth, approval, seat, resignation, vishesh, vyakti 0.6
+movies, movie, download, hollywood, tamil, dubbed, telugu, hd 0.57
+shayari, attitude, love, dil, sad, status, romantic, nahin, mohabbat, girlfriend 0.56
+ganga, festival, dussehra, jayanti, ekadashi, sav, janmashtami, bihu, raksha, dashera 0.55
+computer, processor, tar, virus, system, software, year, input, output 0.54
+yatan, thal, places, tourist, garan, yatak, kuchinda, jaipur, tourism, barot 0.54
+course, bba, llb, pharma, bsc, ba, bachelor, mba, ed, entrance 0.52
+ration, vein, dhaman, artery, card, shan, list, epds, fcs, nfsa 0.51
+jio, phone, tune, caller, recharge, sim, plan, call, reliance, jiotv 0.5
+youtube, channel, video, youtuber, sponsorship, shorts, subscriber, videos, views 0.5
+train, railway, railways, trains, station, vay 0.46
+freelancing, money, kam, earn, paise, online, kamaye, fiverr, earning, freelancer 0.45
+weight, creatine, loss, pathri, at, diet, ghat, patal 0.44
+photo, lightroom, background, image, photoshop, edit, presets, kag, preset, jpeg 0.43
+shram, card, eshram, ram, check, payment, uan, shramik, otp, gov 0.42
+election, nasbah, assembly, bjp, president, elections, powers, veto, gujarat 0.4
+haven, matching, suppliers, supplier, verified, let, found, alibaba, equipment 0.39
+stotram, lord, shiva, shani, dev, ganesha, gan, bhagav, kath 0.39
+kisan, pm, beneficiary, farmer, pm kisan, corner, kyc, status, nidhi, samman 0.38
+Table 14: BERTopic modeling on the Hindi subset of mOSCAR.
+A.4 Heuristics to increase the quality of documents
 We use a set of heuristics to improve the quality of the documents by discarding some text nodes.
@@
 Following previous steps, we keep the text node if it is superior to 5 bytes and we keep the final document if it is superior to 100 bytes.
-A.3 Examples of documents
 Figure 5: Example of a French document.
@@
 Figure 10: Example of an Urdu document.
-A.4 Text-Image similarity and DOM Tree
+A.5 Text-Image similarity and DOM Tree
 As we rely on the DOM Tree to build the documents and the order of appearance of the nodes could differ from HTML rendering, we attempt to assess to what extent it is a relevant way of constructing a multimodal document.
@@
 Figure 11: Relative positions of most relevant images and text nodes with respect to the other modality.
-A.5 Implementation details
-A.5.1 Text deduplication parameters
+A.6 Implementation details
+A.6.1 Text deduplication parameters
 Following previous work, we near-deduplicate documents using MinHashLSH.
@@
 We then compute MinHashes from those vectors with 256 permutations and we finally run Locality Sensitive Hashing with a threshold Jaccard Similarity of 0.8 for finding near-duplicates.
-A.5.2 Training implementation details
+A.6.2 Removing Personal Identifiable Information
+We used regular expressions to detect and remove PII in documents. More precisely, we used:
+email address:
+[…]
+phone number:
+[…]
+[…]
+[…]
+credit card number:
+^(?:4[0-9]{12}(?:[0-9]{3})?|5[1-5]
+[0-9]{14}|3[47][0-9]{13}|3(?:0[0-5]|
+[68][0-9])[0-9]{11}|6(?:011|5[0-9]
+{2})[0-9]{12}|(?:2131|1800|35
+[…]
+IP address:
+^(?:25[0-5]|2[0-4][0-9]|1[0-9]{2}|
+[…]
+[…]
+(?:25[0-5]|2[0-4][0-9]|1[0-9]{2}|
+[…]
+[…]
+passport number: ^[A-Z0-9]{6,15}\$
+For images, we detect faces in the images and distribute the bounding boxes coordinates.
+More precisely, all the images are resized to have a maximum of width and height of 256, keeping aspect ratio.
+The bounding boxes coordinates are therefore computed given this image size but can be extrapolated if images are downloaded in a higher resolution.
+A.6.3 Training implementation details
 We train multilingual OpenFlamingo on mOSCAR and multilingual text-image pairs.
@@
 We use 4 accumulation gradient steps to have an effective batch of size 256 for mOSCAR and 512 for captioning data.
-We train the model on 35M documents and 70M image-text pairs on 8 Nvidia A100 for 120h.
-A.5.3 Evaluation details
+We train the model on 50M documents and 100M image-text pairs on 8 Nvidia A100 for 170h.
+A.6.4 Evaluation details
 Metric #examples Languages
@@
 For captioning tasks, we use the prompt:
-<image>Output:[Caption]<|endofchunk|><image>Output:”,
+<image>Output:[Caption]<|endofchunk|>
+<image>Output:”,
 where [Caption] is replaced by the caption.
 For visual question answering tasks, we use the prompt:
-<image>Question: [Question] Short Answer: [Answer] <|endofchunk|><image>Question: [Question] Short Answer:”,
-where [Question] and [Answer] are replaced by the question and the answer respectively.
+<image>Question: [Question]
+Short Answer: [Answer] <|endofchunk|>
+<image>Question: [Question]
+Short Answer:”,
+where [Question] and [Answer] are replaced by the questions and the answer respectively.
 For multimodal machine translation tasks, we use the prompt:
-<image>Sentence:‘[Caption]’. Translation: [Translation] <|endofchunk|><image>Output:”,
-where [Caption] is replaced by the sentence to translate and [Translation] is replaced by its translation.
+<image>Sentence:‘[Caption]’.
+Translation: [Translation] <|endofchunk|>
+<image>Sentence:‘[Caption]’
+Translation:”,
+where [Caption] is replaced by the sentences to translate and [Translation] is replaced by its translation.
 For MaRVL, we use the prompt:
@@
 For XVNLI, we use the prompt:
-<image> ‘[Statement1]’ - ‘[Statement2]’. entailment, neutral or contradiction? Output: [Answer]<|endofchunk|><image> ‘[Statement1]’ - ‘[Statement2]’. entailment, neutral or contradiction? Output:”,
+<image> ‘[Statement1]’ - ‘[Statement2]’.
+entailment, neutral or contradiction?
+Output: [Answer]<|endofchunk|>
+<image> ‘[Statement1]’ - ‘[Statement2]’.
+entailment, neutral or contradiction?
+Output:”,
 where [Statement1], [Statement2] and [Answer] are replaced by XVNLI test data.
-A.6 Detailed results
+A.7 Detailed results
+Tables 16, 17, 18, 19, 21, 21, 21, 23 and 23 show the detailed results for all languages in which it can be observed that the model trained on mOSCAR outperforms the model trained on captions only by a large margin.
 De En Es Id Ja Ru Tr Zh
 #shots
-0 28.87 33.42 16.35 33.26 04.33 13.99 03.68 18.65
-Multilingual OF 4 53.99 46.53 35.87 48.85 12.44 31.88 11.35 33.64
-mOSCAR + caps. 8 56.05 52.99 37.28 52.51 15.13 32.76 11.69 35.78
-16 55.44 56.78 37.72 49.06 17.10 35.43 12.72 36.82
+0 26.93 29.64 14.07 32.04 02.87 18.07 04.23 07.40
+Multilingual OF 4 54.38 51.47 37.32 47.22 11.06 32.23 13.03 31.71
+mOSCAR + caps. 8 55.09 56.75 34.99 51.60 15.03 34.17 13.63 33.90
+16 61.59 59.89 39.46 51.50 19.63 34.94 14.19 34.49
 0 16.72 24.57 03.80 10.82 02.82 08.20 02.79 06.82
@@
 #shots
-0 06.18 02.02 11.06 09.12 00.90 43.49 18.01 10.41 01.30 20.61 03.21
-Multi. OF 4 21.55 05.93 28.21 21.97 03.38 74.71 36.93 26.27 06.48 39.41 10.34
-full 8 23.15 06.23 31.33 24.53 03.47 75.29 37.58 28.84 07.65 43.40 10.83
-16 23.80 06.86 31.92 24.92 03.59 76.24 38.39 26.78 07.61 43.26 11.78
+0 04.83 02.50 08.52 08.16 00.76 42.57 16.79 12.49 01.26 14.76 03.76
+Multi. OF 4 22.74 06.42 33.73 24.29 02.32 77.98 37.81 31.94 06.78 39.79 15.51
+full 8 22.91 07.41 35.23 25.79 02.95 77.64 38.41 35.46 07.92 42.81 15.85
+16 23.47 08.14 35.96 25.47 02.58 78.18 39.18 31.44 08.42 43.77 16.08
 0 02.24 00.97 06.42 06.46 03.68 10.02 09.32 04.95 01.14 16.15 00.78
@@
 #shots
-0 02.80 01.47 01.85 09.98 11.15 02.07 01.67 18.97 09.63 04.32 15.49
-Multi. OF 4 10.62 07.48 05.51 22.63 27.88 16.87 09.24 42.30 22.35 14.20 29.35
-full 8 11.12 07.54 05.91 25.39 29.34 19.35 09.99 46.79 23.54 15.43 32.69
-16 12.18 07.71 06.03 23.89 29.17 18.84 09.75 46.95 23.79 15.93 31.98
+0 02.79 02.00 01.51 09.96 11.53 00.92 00.58 16.11 08.31 03.94 13.37
+Multi. OF 4 11.03 10.87 05.87 25.88 29.53 17.45 10.85 46.22 25.18 15.36 31.32
+full 8 11.61 12.00 06.91 29.68 29.34 20.13 12.01 47.58 27.08 17.80 33.29
+16 12.74 11.40 07.03 26.73 30.43 20.57 11.07 49.33 27.07 17.15 32.79
 0 02.29 00.97 03.51 02.98 07.96 01.85 01.05 04.88 05.78 00.92 09.79
@@
 #shots
-0 02.19 06.80 11.45 00.87 08.36 03.11 03.04 21.99 07.19
-Multi. OF 4 05.63 20.50 25.59 02.29 21.66 10.87 10.43 38.43 19.34
-full 8 06.04 23.11 26.44 03.06 25.19 12.95 10.56 39.75 20.18
-16 06.43 21.15 26.51 03.42 25.31 13.57 10.45 40.85 20.46
+0 01.84 04.72 11.09 00.88 05.49 02.86 02.08 11.34 03.29
+Multi. OF 4 06.08 21.46 30.24 03.46 23.14 10.75 11.35 32.70 19.57
+full 8 07.10 21.78 30.26 03.76 25.17 12.83 12.26 35.86 20.11
+16 06.95 22.63 32.07 04.52 25.23 13.38 12.29 37.12 20.71
 0 02.24 01.93 04.55 00.67 02.34 02.68 00.80 08.55 02.70
@@
 #shots
-0 20.99 22.57 32.60 22.45 26.44 23.54 25.31 26.77
-Multilingual OF 4 25.72 32.71 38.02 30.74 31.49 31.53 31.75 33.26
-mOSCAR + caps. 8 26.87 35.60 39.16 34.25 32.25 34.93 33.59 34.55
-16 29.11 37.01 40.64 36.37 34.20 37.15 34.87 36.29
+0 22.76 25.72 34.24 26.68 26.89 26.73 25.28 27.32
+Multilingual OF 4 26.72 32.57 37.91 32.54 31.88 32.35 31.28 33.4
+mOSCAR + caps. 8 28.07 35.15 39.44 35.14 32.94 35.59 33.58 34.04
+16 29.64 37.33 40.09 35.55 34.06 36.27 34.50 35.36
 0 10.54 06.51 10.43 07.74 07.50 07.79 08.62 09.84
@@
 #shots
-0 34.24 15.91 18.08 16.43 13.73 25.75 13.36
-Multi. OF 4 36.19 21.21 20.77 19.29 19.01 31.72 19.13
-mOSCAR + caps 8 36.96 17.80 20.00 20.36 17.96 33.96 23.83
-16 36.19 18.18 21.92 20.71 16.90 41.04 22.74
+0 36.58 28.03 20.38 18.21 15.49 24.25 13.36
+Multi. OF 4 38.13 30.03 23.08 21.43 17.61 31.72 22.02
+mOSCAR + caps 8 38.52 29.55 24.62 20.00 17.61 25.27 23.83
+16 35.80 31.82 25.00 23.93 19.01 33.96 22.74
 0 09.73 00.38 07.69 01.43 00.00 05.22 03.61
@@
 Random chance 50.00 50.00 50.00 50.00 50.00
-0 50.00 49.46 49.76 49.83 49.80
-Multilingual OF 4 49.91 49.55 49.28 49.83 49.80
-mOSCAR + caps 8 50.62 48.65 49.68 49.83 49.80
-16 50.18 49.01 49.76 49.83 49.90
+0 50.09 49.46 49.60 49.83 48.81
+Multilingual OF 4 49.91 48.19 49.68 50.42 50.00
+mOSCAR + caps 8 53.55 50.72 49.76 51.78 51.58
+16 48.94 49.82 49.20 50.25 50.99
 0 51.33 49.01 49.52 49.83 49.70
@@
 Random chance 33.33 33.33 33.33 33.33 33.33
-0 32.90 34.02 31.44 33.85 32.82
-Multilingual OF. 4 36.94 36.17 34.19 34.54 38.49
-mOSCAR + caps. 8 37.80 39.86 34.97 35.74 37.46
-16 34.97 38.92 32.30 34.97 35.65
+0 33.51 34.62 33.08 34.02 34.19
+Multilingual OF. 4 33.08 33.59 33.42 34.45 35.82
+mOSCAR + caps. 8 35.91 38.75 35.14 36.08 37.11
+16 34.11 36.60 33.93 34.54 35.05
 0 35.48 34.02 33.51 34.45 31.36
@@
 #shots
-0 3.22 28.86 36.01
-Multi. OF 4 3.16 28.99 36.22
-full 8 3.44 28.76 37.41
-16 3.73 29.19 37.53
+0 2.82 28.45 37.47
+Multi. OF 4 3.12 29.20 37.49
+full 8 3.14 29.62 37.99
+16 3.34 29.41 38.79
 0 0.00 00.00 00.00
@@
 #shots
-0 59.09 63.67 68.51
-Multi. OF 4 58.77 63.67 68.51
-full 8 56.82 63.67 68.51
-16 57.79 62.67 67.86
+0 56.49 65.67 67.86
+Multi. OF 4 57.47 64.00 68.18
+full 8 58.44 64.33 67.86
+16 58.11 62.67 66.23
 0 58.12 61.67 64.29
@@
 […]
+# shots xFlickR&CO XM3600 xGQA MaXM MaRVL XVNLI Multi30k CoMMuTE
+InternVL2 4B 0 16.21 07.02 12.38 06.35 53.14 33.85 26.99 66.93
+4 24.89 09.53 26.05 14.72 54.22 35.72 26.68 64.22
+PaliGemma 3B 0 28.28 24.49 42.68 33.42 51.48 39.36 17.98 62.78
+Idefics2 8B 0 27.11 15.94 22.53 28.99 63.18 50.33 30.19 67.13
+Llava-NeXT 8B 0 23.67 14.70 25.48 15.17 60.50 45.40 29.40 66.37
+Multi. OF 3B (ours) 0 16.91 07.45 26.95 22.23 49.56 33.88 22.91 63.34
+4 34.80 22.18 32.23 26.33 49.64 34.07 23.27 63.22
+Table 24: Results averaged across languages. Bold is best result.
+A.8 Comparison with state-of-the-art mLLMs
+We computed the results for different state-of-the-art models of similar sizes as multilingual Open Flamingo namely: (1) InternVL2-4B [^26] OpenGVLab/InternVL2-4B (2) PaliGemma272727google/paligemma-3b-pt-224 (3) Idefics2-8B [^28] HuggingFaceM4/idefics2-8b and (3) Llava-NeXT 8B292929llava-hf/llama3-llava-next-8b-hf.
+InternVL2 and PaliGemma are trained on multilingual and multimodal data while Llava-NeXT and Idefics2 are trained on English multimodal datasets.
+Table 24 shows results averaged across languages for different state-of-the-art mLLMs of sizes from 3b to 8B.
+These results highlights multiple things: (1) getting results significantly better than random (MaRVL and XVNLI) requires instruction-tuning data as Idefics2 and Llava-NeXT were both trained on instruction-tuning multimodal datasets (2) English-only still gets decent results on multilingual benchmarks despite not having been trained on multilingual and multimodal data, probably due to their underlying LLM being multilingual (3) multilingual Open Flamingo (trained on mOSCAR and captions) gets superior results to InternVL2-4B on VQA benchmarks and captioning benchmarks but inferior to PaliGemma-3B mainly due to the fact that it was trained on much less data and the quality of the captions used to train multilingual Open Flamingo may not be as good as the WebLI dataset used to train PaliGemma.